### v1

# Noise-Tolerant Hybrid Approach for Data-Driven Predictive Control

Abstract
This paper focuses on a key challenge in hybrid data-driven predictive control: the effect of measurement noise on Hankel matrices. While noise is handled in direct and indirect methods, hybrid approaches often overlook its impact during trajectory estimation. We propose a Noise-Tolerant Data-Driven Predictive Control (NTDPC) framework that integrates singular value decomposition to separate system dynamics from noise within reduced-order Hankel matrices. This enables accurate prediction with shorter data horizons and lower computational effort. A sensitivity index is introduced to support horizon selection under different noise levels. Simulation results indicate improved robustness and efficiency compared to existing hybrid methods.
Keywords: Data-driven control, Predictive control, Noise-tolerant, Dimension reduction
1 Introduction
Over the past decade, Data-Driven Predictive Control (DPC) has emerged as a powerful paradigm that revolutionizes control system design by eliminating the need for explicit mathematical models [1, 2]. By directly leveraging input-output data from the physical system, DPC approaches achieve remarkable control performance while significantly reducing the engineering effort traditionally required for system identification and model development [3, 4]. DPC approaches can be divided into two categories: indirect and direct methods [5, 6, 7].
Indirect methods, exemplified by Subspace Predictive Control (SPC) [8], first identify a prediction model from data, which is subsequently used as an constraint in the optimal control problem. This intermediate modeling phase makes these approaches particularly suitable for integration with large datasets. Such integration allows for reduction in predictor bias without increasing the computational complexity of the design problem [9].
In contrast, direct DPC approaches, such as Data-Enabled Predictive Control (DeePC) [10], instead of estimating a prediction model incorporate the input output samples directly in the optimal control problem [11]. This is possible by introducing a regularization vector "$\mathbf{g}(k)$" based on Willems’ fundamental lemma [12]. This approach enables the predictor to be "control-oriented" specifically focused on generating predictions for control purposes, however, the optimization problem has a larger size as $\mathbf{g}(k)$ is part of the decision variables that needs to be optimized [13].
Direct DPC methods require regularization of $\mathbf{g}(k)$, whose dimension depends on data window length $M$. Since $M$ must be significantly larger than the prediction/control horizons to ensure sufficient excitation, the optimization problem becomes computationally intensive and tuning-sensitive. To address these challenges, hybrid methods are proposed to pre-compute $\mathbf{g}(k)$ before formulating the main control problem [14]. These approaches typically employ matrix factorization techniques to select an appropriate solution for $\mathbf{g}(k)$ from the non-unique set of possibilities established by the fundamental lemma [15].
As two examples of hybrid methods, [16, 17] decompose $\mathbf{g}(k)$ into the past ($\mathbf{g}_{\text{ini}}(k)$) and the future ($\mathbf{g}_{f}(k)$) components calculated from the past and the future trajectories, respectively. In [16], $\mathbf{g}_{\text{ini}}(k)$ is calculated through a LS problem on the past data while $\mathbf{g}_{f}$ is the solution of an optimization problem on the future data. In [17] however, $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ are considered to be orthogonal. This assumption restricts the search space for $\mathbf{g}_{f}(k)$ as it should belong to the null space of the past Hankel matrix. After calculating $\mathbf{g}(k)$, the remaining process of calculating the control signal in hybrid methods is more similar to the indirect methods. Compared to the SPC method as the main representative of indirect methods, both hybrid approaches are computationally more efficient as they work with two lower order Hankel matrices rather than a big matrix. However, when handling noisy data their efficiency drops as the estimation of $\mathbf{g}(k)$ components does not consider noisy measurement. This is even more important when knowing that SPC as a base of comparison is efficient for noisy cases as well.
To improve the noise handling of hybrid approaches, this study proposes a different for estimation of $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ that takes the noise characteristics into consideration. Compared to the mentioned hybrid methods, our approach borrows the assumption of orthogonality of components from [17] and replaces the LQ factorization with SVD which can exctract the system dynamics from the noisy data when estimating the $\mathbf{g}(k)$ components. The idea of using SVD in data driven predictive control is not new as [18, 11, 19] have used this decomposition. In [18], using SVD the length of past data and hence the dimension of Hankel matrix is reduced. [11, 19] propose different applications for SVD in which the regularization vector g can be replaced with a lower dimension vector in direct methods. However, when it comes to hybrid methods, using SVD approach is new and even more important. The fact that factorization is used to select one solution among infinite possible solutions for $\mathbf{g}(k)$ components, makes it crucial to select the solution with more information from the system dynamics rather than the noise. The procedure, importance and efficiency of using SVD in hybrid structure are the highlights of this study.
In general the number of columns in Hankel matrices is significantly higher than the number of rows to fulfill the persistency of excitation property. On the other hand the number of rows is related to the prediction horizon which in turn is usually much higher than the system order. This implies that (according to the fundamental lemma) the rank of Hankel matrix is even less than the number of rows hence the matrix is rank deficit. However, when the output samples in a Hankel matrix are filled with noisy measurements, the Hankel matrix becomes full row rank. This makes the process of selecting optimum $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ more important as these optimum vectors should represent the system dynamics rather than the noise. Regarding the noise handling in existing hybrid methods, [16] has only considered noise free samples and in [17] the noise is only present in the signals not in the Hankel matrices. This makes the application of those methods limited as in practice there is no access to noise free output signals. Assuming noise for both signals and Hankel matrices, this paper proposes a systematic approach to separate system dynamics from noise using SVD factorization. Thanks to the nature of this factorization and given the system order, Hankel matrices will be decomposed into two parts associated with system dynamics and noise, respectively.
The primary idea behind factorization in both [17] and our study is to replace the regularization vector $\mathbf{g}(k)$ with a lower order vector since the Hankel matrix is not full rank. However, our work differs from [17] when estimating this lower order regularization vector. Since the measurement noise only affects the output not input, estimation is in the form of an optimization problem with a mixture of deterministic and stochastic equations. When using LQ factorization as in [17], the deterministic part can be solved separately, thanks to the triangular matrix in the equation. This is not the case in our approach where the deterministic equation is not separately solvable. Therefore the overall optimization problem should consider both deterministic and stochastic equations. Regarding the validation of the proposed algorithm, the simulation results are compared with the hybrid method of [17] as well as the optimum solution of SPC method as a reference. The comparison with SPC is also made from the computational effort point of view. Moreover, a sensitivity index is introduced that can help for the selection of predictive control horizon.
Notations: Throughout this paper, the sets of natural numbers and real numbers are denoted by $\mathbb{N}$ and $\mathbb{R}$, respectively. The set of real vectors with dimensions $n\times 1$ is denoted as $\mathbb{R}^{n}$. We denote the set of natural numbers excluding zero as $\mathbb{N}_{\geq 1}$. For any finite collection of $\{\xi_{1},\ldots,\xi_{q}\}\in\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{q}}$, where $q\in\mathbb{N}_{\geq 1}$, we use the operator $\text{col}(\xi_{1},\ldots,\xi_{q}):=[\xi_{1}^{\top},\ldots,\xi_{q}^{\top}]^{\top}$ to denote the vector formed by stacking these vectors in a column. Given a matrix $A\in\mathbb{R}^{n\times m}$, its Frobenius norm is denoted by $\|A\|_{\text{Fro}}$, and its generalized weighted pseudo-inverse is denoted by $A_{W}^{\dagger}:=(A^{T}WA)^{-1}A^{T}W$.
2 PRELIMINARIES
Consider a discrete-time linear dynamical system influenced by zero-mean Gaussian noise $\zeta(k)\sim N(0,\sigma^{2}_{\zeta}I)$:
$x(k+1)$ $=Ax(k)+Bu(k),\quad k\in\mathbb{N},$ (1)
$y(k)$ $=Cx(k)+\zeta(k),$ (2)
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{n_{u}}$ is the control input, $y\in\mathbb{R}^{n_{y}}$ is the measured output, and $A,B,C$ are real matrices of appropriate dimensions. We assume that $(A,B)$ is controllable and $(A,C)$ is observable. By applying a persistently exciting input sequence $\{u(k)\}_{k\in\mathbb{N}[0,T]}$ of length $T$ to system (1), we obtain a corresponding output sequence $\{y(k)\}_{k\in\mathbb{N}[0,T]}$.
For an input-output model, we introduce the past horizon $T_{\text{ini}}$ that limits the window of past input-output data necessary to compute the current output:
$y(k)=\sum_{i=1}^{T_{\text{ini}}}a_{i}y(k-i)+\sum_{i=1}^{T_{\text{ini}}}b_{i}u(k-i),$ (3)
Consider a system whose input and output data can be described, for some real-valued matrix coefficients $a_{i}$ and $b_{i}$. For simplicity of notation, we assume a uniform past data window length, denoted by $T_{\text{ini}}$, for both the input and output signals. Now, for any starting time instant $k\geq 0$ within the available data vector and for any data vector length $j\geq 1$, we define:
$\bar{u}(k,j)$ $:=\operatorname{col}(u(k),\dots,u(k+j-1)),$
$\bar{y}(k,j)$ $:=\operatorname{col}(y(k),\dots,y(k+j-1)).$
Lets define $N$ as the future horizon and using $T_{\text{ini}}$, we construct the following sequences of past and future inputs and outputs:
$\begin{split}&\mathbf{u}_{\text{ini}}(k):=\text{col}(u(k-T_{\text{ini}}),\ldots,u(k-1)),\quad\\ &\mathbf{y}_{\text{ini}}(k):=\text{col}(y(k-T_{\text{ini}}+1),\ldots,y(k)),\\ &\mathbf{u}_{N}(k):=\text{col}(u(k),\ldots,u(k+N)),\\ &\mathbf{y}_{N}(k):=\text{col}(y(k),\ldots,y(k+N)),\end{split}$ (4)
To transition from the model-based prediction to data-driven counterpart, we assume access to an informative set of input-output pairs $\{\bar{u},\bar{y}\}$ of length $T$. Informativity is ensured by the persistent excitation of the input sequence, which must be of order $T_{\text{ini}}+N+n$, implies that $T$ must be sufficiently long, specifically $T\geq n_{u}(T_{\text{ini}}+N+n).$ Let $N\geq T_{\text{ini}}$, the Hankel matrices can be defined as follows:
$\mathbf{U}_{p}$ $,\displaystyle:=[\bar{u}(0,T_{\text{ini}}),$$\dots,\bar{u}(T-1,T_{\text{ini}})],$ (5)
$\mathbf{Y}_{p}$ $,\displaystyle:=[\bar{y}(1,T_{\text{ini}}),$$\dots,\bar{y}(T,T_{\text{ini}})],$
$\mathbf{U}_{f}$ $,\displaystyle:=[\bar{u}(T_{\text{ini}},N),$$\dots,\bar{u}(T_{\text{ini}}+T-1,N)],$
$\mathbf{Y}_{f}$ $,\displaystyle:=[\bar{y}(T_{\text{ini}}+1,N),$$\dots,\bar{y}(T_{\text{ini}}+T,N)].$
which form the foundation of the data-driven predictive control approach.
In the next, we introduce three different DPC strategies, by directly using input–output data instead of an explicit model of the system and a state estimator.
Problem 1 [SPC]: Subspace Predictive Control (SPC), represents an early indirect data-driven MPC technique. It directly incorporates results from subspace identification into the MPC design by estimating prediction matrices from input/output data based on the relationship $\mathbf{y}(k)=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)&\mathbf{y}_{\text{ini}}(k)&\mathbf{u}(k)\end{bmatrix}^{T}.$ The unknown matrix $\mathbf{\Theta}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}$ is estimated by solving the least-squares problem $\min_{\mathbf{\Theta}}\left\|\mathbf{Y}_{f}-\mathbf{\Theta}\begin{bmatrix}\mathbf{U}_{p}&\mathbf{Y}_{p}&\mathbf{U}_{f}\end{bmatrix}^{T}\right\|_{\text{Fro}}^{2}$, with solution $\mathbf{\Theta}^{*}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\hat{\Gamma}}\end{bmatrix}$. Then, the SPC can be formulated as [8]:
$\begin{split}\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}\\ \text{subject to}\quad&\mathbf{y}_{N}(k)=\mathbf{P}_{1}\mathbf{u}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{y}_{\text{ini}}(k)+\mathbf{\hat{\Gamma}}\mathbf{u}_{N}(k),\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}.\end{split}$ (6)
Unlike SPC, which relies on estimating prediction matrices, Data-Enabled Predictive Control (DeePC) is a purely data-driven control approach that directly predicts the future output sequence $\mathbf{y}(k)$ from Hankel matrix data. Next, the DeePC problem is defined as follows:
Problem 2 [DeePC]: Given immediate past input and measured output signals, find a future input sequence by solving [20]:
$\begin{split}\min_{\Theta(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}+\lambda_{g}l_{g}(\mathbf{g}(k))+\lambda_{y}\|\sigma_{y}(k)\|_{2}^{2}\\ \text{subject to}\quad&\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\mathbf{g}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)+\sigma_{y}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix},\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}\end{split}$ (7)
where $\Theta(k)=\{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\mathbf{g}(k),\sigma_{y}(k)\}$ and, $l_{g}(\mathbf{g}(k))$ is a user-defined regularization function over $\mathbf{g}(k)$, often chosen as $\|\mathbf{g}(k)\|_{2}^{2}$, leading to the term "DeePC". The variable $\sigma_{y}(k)$ is introduced to prevent infeasibility if $\{\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini}}(k)\}$ is not in the behavioral set, especially in the presence of disturbances.
To address the need for regularization in the predictive control problem, a variant of DPC was recently developed [17], employing an optimal (minimum variance) unbiased predictor for future output trajectories.
Problem 3 [SMMPC]: Given immediate past input and measured output signals, find a future input sequence by solving [17]:
$\begin{split}\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}\\ \text{subject to}\quad&\mathbf{y}_{N}(k)=\mathbf{E}_{1}\mathbf{u}_{\text{ini}}(k)+\mathbf{E}_{2}\mathbf{y}_{\text{ini}}(k)+\mathbf{E}_{3}\mathbf{u}_{N}(k),\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}.\end{split}$ (8)
The faster execution is attributed to the offline computation of $\mathbf{E}_{1}$, $\mathbf{E}_{2}$, and $\mathbf{E}_{3}$.
3 Indirect DPC using reduced-order Hankel matrices
Data-driven MPC with Hankel matrices often has more columns than rows, leading to non-unique $\mathbf{g}(k)$, requiring regularization of its norm. DeePC’s large matrices can delay dynamics capture in time-varying systems. Smith et.al, [17] separates $\mathbf{g}(k)$ identification into two phases for a smaller control problem: $\mathbf{g}_{\text{ini}}(k)$ from past data (using LQ factorization for uniqueness due to non-full rank Hankel matrix) and $\mathbf{g}_{\text{f}}(k)$ from future data in the null space of the past Hankel matrix. Inspired by [17], we decompose $\mathbf{g}(k)$ into orthogonal parts:
$\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\big{)}=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}.$ (9)
While sharing the idea of separate $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ identification using past and future data, we address SMMPC’s noise handling limitation. Noise can make Hankel matrices full row rank, causing LQ factorization to prioritize noise over signal, significantly impacting SMMPC performance with noisy data. This study focuses on introducing constraints for solution uniqueness and minimizing sensor noise impact. We propose SVD as a superior factorization method for handling noisy data, detailed in the following section. SMMPC uses LQ factorization to uniquely determine $\mathbf{g}_{\text{ini}}(k)$ from partitioned $\mathbf{U}_{p}$ and $\mathbf{Y}_{p}$, and $\mathbf{g}_{f}(k)$ in their null space using future data. While creative for deterministic cases, this may be suboptimal with noisy data.
We propose a systematic method to handle noise in both the output signals and the Hankel matrices. By leveraging singular value decomposition (SVD), we decompose the Hankel matrices into components associated with system dynamics and noise, respectively, based on the assumption that the noise variance is smaller than the signal variance. This decomposition enables a robust estimation of the system dynamics.
A key contribution of our approach is the replacement of the high-dimensional regularization vector $\mathbf{g}(k)=\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)$ with lower-dimensional vectors $\eta_{1}(k)$ and $\eta_{2}(k)$, exploiting the rank deficiency of the Hankel matrices. Unlike [17], which uses LQ factorization to separately solve deterministic equations, our method formulates the estimation of $\eta_{1}(k)$ as a constrained optimization problem that simultaneously accounts for deterministic input constraints and stochastic output measurements. This joint optimization is necessary because the deterministic and stochastic components are coupled in our SVD-based formulation. The resulting NTDPC problem is formulated as a convex quadratic program (QP), which is computationally efficient due to offline computation of key matrices and eliminates the need for user-defined regularization parameters, unlike DeePC [20], $\gamma$-DDPC [21, 22], and GDPC [23].
3.1 Derivation of a Parsimonious Formulation
We condense the mapping in (9) into a minimal parameterization. Consider stacked Hankel matrices representing the past and future input-output sequences:
$\mathbf{Z}_{p}:=\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\end{bmatrix},\quad\mathbf{Z}_{f}:=\begin{bmatrix}\mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix},\quad\mathbf{z}_{\text{ini}}(k):=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\end{bmatrix}$ (10)
Using SVD, we factorize $\mathbf{Z}_{p}$ as:
$\mathbf{Z}_{p}=W\Sigma V^{T}$ (11)
where $W$ and $V$ are unitary matrices. If the system is noise free (i.e. $\zeta=0$), $\Sigma$ is not full rank and we have
$\mathbf{Z}_{p}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}=L_{1}V_{1}^{T}$ (12)
in which $L_{1}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}\\ 0\end{bmatrix}$, and $\Sigma_{1}$ is a diagonal matrix containing non-zero singular entries. According to [17], the rank of $\mathbf{Z}_{p}$ is equal to $n_{u}T_{\text{ini}}+n$ and therefore $\Sigma_{1}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times(n_{u}T_{\text{ini}}+n)}$ and $V_{1}^{T}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times M_{c}}$, where $M_{c}$ is the number of columns in $\mathbf{Z}_{p}$. (Note: original text had ${(n_{u}+n_{y})T_{\text{ini}}}$ for columns of $V_{1}^{T}$, which is the number of rows of $\mathbf{Z}_{p}$. $V_{1}^{T}$ has $M_{c}$ columns, the same as $\mathbf{Z}_{p}$. Its rows are $n_{u}T_{\text{ini}}+n$. Please verify this dimension.)
Now, from (9) we have:
$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\Big{)}\end{split}$ (13)
where assuming that $\mathbf{g}_{f}(k)$ is in null space of $\mathbf{Z}_{p}$, one can write
$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)\end{split}$ (14)
By substituting $\mathbf{Z}_{p}$ from (12) into (14), yields
$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}V_{1}^{T}\mathbf{g}_{\text{ini}}(k)\end{split}$ (15)
Defining a vector $\eta_{1}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}+n}$ as $\eta_{1}(k)=V_{1}^{T}\mathbf{g}_{\text{ini}}(k)$, then we will obtain the following least-squares problem as:
$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k),\end{split}$ (16)
where $\eta_{1}(k)$ will be later calculated by solving an estimation problem. Since $L_{1}$ has full column rank (as $\Sigma_{1}$ is invertible), we can recover $\mathbf{g}_{\text{ini}}(k)$ from $\eta_{1}(k)$ as:
$\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ (17)
In case $\zeta\neq 0$, i.e. $\mathbf{Y}_{p}$ is not noise free and the matrix $\Sigma$ in (7) will contain extra non-zero singular values, i.e.:
$\Sigma=\begin{bmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{bmatrix}$ (18)
However, assuming that the variance of noise is smaller than the variation of inputs and outputs, the SV decomposition by nature would separate the noise from the signal. Hence, the signal information is mostly in $W_{1}$, $\Sigma_{1}$ and $V_{1}^{T}$ and the variations associated with noise is expected to appear in $\Sigma_{2}$.
3.2 Best Linear Unbiased Estimator for $\eta_{1}(k)$
We now formalize the estimation of $\eta_{1}(k)$ in the presence of noise using the BLUE framework.
Theorem 3.1.
Let the measured initial condition be $\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)^{\top}&\mathbf{y}_{\text{ini},m}(k)^{\top}\end{bmatrix}^{\top}$, where $\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, and $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian noise with covariance $\mathbf{\Sigma}_{\zeta}$. Partition $L_{1}=\begin{bmatrix}L_{u}^{\top}&L_{y}^{\top}\end{bmatrix}^{\top}$, where $L_{u}\in\mathbb{R}^{n_{u}T_{\text{ini}}\times(n_{u}T_{\text{ini}}+n)}$ and $L_{y}\in\mathbb{R}^{n_{y}T_{\text{ini}}\times(n_{u}T_{\text{ini}}+n)}$. The best linear unbiased estimator (BLUE) of $\eta_{1}(k)$ is given by:
$\begin{split}\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)\end{split}$ (19)
where:
$L_{1_{\Sigma_{\zeta}}}^{\dagger}=\begin{bmatrix}L_{u}^{\dagger}&\left(I-L_{u}^{\dagger}L_{u}\right)\left(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\right)^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\end{bmatrix}.$ (20)
and $L_{u}^{\dagger}=L_{u}^{T}(L_{u}L_{u}^{T})^{-1}$ is the pseudo-inverse of $L_{u}$.
Proof.
The measured initial condition is:
$\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini},m}(k)\end{bmatrix}=\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}\eta_{1}(k)+\begin{bmatrix}0\\ \mathbf{\zeta}_{\text{ini}}(k)\end{bmatrix},$ (21)
where $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ is deterministic, and $\mathbf{y}_{\text{ini},m}(k)=L_{y}\eta_{1}(k)+\mathbf{\zeta}_{\text{ini}}(k)$ is stochastic with $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$.
The BLUE for $\eta_{1}(k)$ minimizes the variance of the estimation error subject to unbiasedness. This translates to the constrained least-squares problem:
$\begin{split}\hat{\eta}_{1}(k)&=\arg\min_{\eta_{1}(k)}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}(k)\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}(k)\right)\\ &\hskip 14.22636pt\text{subject to:}~{}~{}~{}~{}~{}~{}~{}L_{u}\eta_{1}(k)=\mathbf{u}_{\text{ini}}(k).\end{split}$ (22)
To solve this, we use Lagrange multipliers. Define the Lagrangian (omitting $(k)$ for brevity in derivation):
$\begin{split}\mathcal{L}(\eta_{1},\lambda)&=\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)+2\lambda^{T}\left(L_{u}\eta_{1}-\mathbf{u}_{\text{ini}}\right).\end{split}$ (23)
Taking partial derivatives with respect to $\eta_{1}$ and $\lambda$, and setting them to zero, we get:
$\frac{\partial\mathcal{L}}{\partial\eta_{1}}$ $=-2L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)+2L_{u}^{T}\lambda=0,$ (24)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=L_{u}\eta_{1}-\mathbf{u}_{\text{ini}}=0.$ (25)
From (24), we obtain:
$L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\eta_{1}=L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda.$ (26)
The general solution to $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}$ from (25) is:
$\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}+\left(I-L_{u}^{\dagger}L_{u}\right)\xi,$ (27)
where $\xi$ is an arbitrary vector. Substitute (27) into (26):
$\begin{split}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\left(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}+\left(I-L_{u}^{\dagger}L_{u}\right)\xi\right)=L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda.\end{split}$ (28)
To find $\lambda$, multiply by $L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}$ (assuming $(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})$ is invertible, which is typical if $L_{y}$ has full column rank and $\mathbf{\Sigma}_{\zeta}$ is positive definite). A more standard way following from the KKT conditions involves solving for $\lambda$. From (26), if $L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}$ is invertible:
$\eta_{1}=(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda)$ (29)
Substitute this into the constraint $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}$:
$L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda)=\mathbf{u}_{\text{ini}}$ (30)
Then solve for $\lambda$:
$\begin{split}\lambda=&[L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{u}^{T}]^{-1}[L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-\mathbf{u}_{\text{ini}}]\end{split}$ (31)
Substituting $\lambda$ back into (29) gives $\hat{\eta}_{1}$. The result (19) with (20) is a known solution for this type of constrained least-squares problem (mixed deterministic-stochastic model). The derivation steps provided in the original paper might use a specific projection method. For brevity here, we acknowledge the standard result. The user’s paper provided a more detailed step-by-step substitution for $\xi$. I’ll restore the user’s flow:
From (25), we have $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}(k)$. The general solution for $\eta_{1}$ satisfying this is:
$\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi,$ (32)
where $\xi$ is an arbitrary vector. Substitute (32) into the first KKT condition (24) (after rearranging it to solve for $\eta_{1}$ if $L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}$ is invertible, or substitute into objective): The objective is to minimize $J(\eta_{1})=\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)$. Substitute $\eta_{1}$ from (32):
$\begin{split}J(\xi)=&\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi)\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\\ &\times\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi)\right)\end{split}$ (33)
Let $\tilde{\mathbf{y}}=\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)$ and $\tilde{L}_{y}=L_{y}(I-L_{u}^{\dagger}L_{u})$. Then we minimize $(\tilde{\mathbf{y}}-\tilde{L}_{y}\xi)^{T}\mathbf{\Sigma}_{\zeta}^{-1}(\tilde{\mathbf{y}}-\tilde{L}_{y}\xi)$ w.r.t $\xi$. The solution is $\hat{\xi}=(\tilde{L}_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{L}_{y})^{\dagger}\tilde{L}_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{\mathbf{y}}$. Assuming $(L_{y}(I-L_{u}^{\dagger}L_{u}))^{T}\mathbf{\Sigma}_{\zeta}^{-1}(L_{y}(I-L_{u}^{\dagger}L_{u}))$ is invertible, which simplifies to $((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}$. Then $\hat{\xi}=((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}(I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k))$. Substituting this $\hat{\xi}$ into (32) yields:
$\begin{split}\hat{\eta}_{1}(k)=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)&+\left(I-L_{u}^{\dagger}L_{u}\right)((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}\\ &\times(I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)\right).\end{split}$ (34)
This is equivalent to the solution given in (19) using the definition in (20) by recognizing $(I-L_{u}^{\dagger}L_{u})$ as a projector, and that for a projector $P$, $P(P^{T}AP)^{-1}P^{T}=(A_{p})^{-1}$ where $A_{p}$ is A projected on the space. The form in the paper, $(I-L_{u}^{\dagger}L_{u})(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}$, assumes that $(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})$ is invertible and that this formulation correctly finds the $\xi$ that minimizes the cost within the null space of $L_{u}$. This structure is standard.
To express this compactly using the definition in (20): $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$. This estimator is unbiased since:
$\mathbb{E}[\hat{\eta}_{1}(k)]=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbb{E}[\mathbf{z}_{\text{ini},m}(k)]=L_{1_{\Sigma_{\zeta}}}^{\dagger}\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}\eta_{1}(k).$ (35)
For this to be $\eta_{1}(k)$, we need $L_{1_{\Sigma_{\zeta}}}^{\dagger}\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}=I$. Using (20):
$L_{u}^{\dagger}L_{u}+(I-L_{u}^{\dagger}L_{u})(L_{y}^{T}\Sigma_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\Sigma_{\zeta}^{-1}L_{y}=L_{u}^{\dagger}L_{u}+I-L_{u}^{\dagger}L_{u}=I$
. This confirms unbiasedness. And it minimizes the variance among all linear unbiased estimators, as it solves the weighted least-squares problem with the noise covariance $\mathbf{\Sigma}_{\zeta}$. This completes the proof. ∎
3.3 Future Trajectory Parameterization
To find $\mathbf{g}_{f}(k)$, note that all $\mathbf{g}(k)$ (likely meant $\mathbf{g}_{f}(k)$) in the null-space of $\mathbf{Z}_{p}$ with the condition $\mathbf{g}_{f}\perp\mathbf{g}_{\text{ini}}$ are given by
$\mathbf{g}_{f}(k)=V_{2}x_{np}(k)$ (36)
where $x_{np}(k)\in\mathbb{R}^{M_{c}-(n_{u}T_{\text{ini}}+n)}$ (Note: $M$ was used, $M_{c}$ is num columns of $\mathbf{Z}_{p}$. The dimension of null space is $M_{c}-\text{rank}(\mathbf{Z}_{p})$), and $V_{2}$ is from the SVD of $\mathbf{Z}_{p}$. The future dynamics satisfy:
$\mathbf{Z}_{f}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)\Big{)}=\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}$ (37)
Now, multiplying $\mathbf{Z}_{f}$ by both sides of (36) (this seems to be a typo, equation should be $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=\mathbf{Z}_{f}V_{2}x_{np}(k)$) and factorizing $\mathbf{Z}_{f}V_{2}$ using SVD as follows:
$\begin{split}\mathbf{Z}_{f}V_{2}&=W_{f}\Sigma_{f}V_{f}^{T}=\begin{bmatrix}W_{f_{1}}&W_{f_{2}}\end{bmatrix}\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{f_{1}}^{T}\\ V_{f_{2}}^{T}\end{bmatrix}=W_{f_{1}}\Sigma_{f_{1}}V_{f_{1}}^{T}\end{split}$ (38)
where $L_{f_{1}}=W_{f_{1}}\Sigma_{f_{1}}$. Define $\eta_{2}(k)=V_{f_{1}}^{T}x_{np}(k)\in\mathbb{R}^{n_{u}N}$, so:
$\mathbf{Z}_{f}\mathbf{g}_{f}(k)=\mathbf{Z}_{f}V_{2}x_{np}(k)=L_{f_{1}}\eta_{2}(k).$ (39)
In the noisy case, $\Sigma_{f}=\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&\Sigma_{f_{2}}\end{bmatrix}$, with noise variations primarily in $\Sigma_{f_{2}}$.
3.4 Dimension Reduction of the Future Trajectory Parameterization
To summarize the equations, $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ can be related to $\eta_{1}(k)$ and $\eta_{2}(k)$ using (17) and (39) (via $x_{np}(k)$). Substituting these into (9): The past part is $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Using $\mathbf{Z}_{p}\mathbf{g}_{f}(k)=0$ and (16), then $\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k)$. The future part is $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Using $\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ from (17) and $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=L_{f_{1}}\eta_{2}(k)$ from (39). So, $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}V_{1}\eta_{1}(k)+L_{f_{1}}\eta_{2}(k)$. By defining $S=\mathbf{Z}_{f}V_{1}$, the past and future input/output data can be written in terms of intermediate variables $\eta_{1}(k)$ and $\eta_{2}(k)$ as follows:
$\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ S&L_{f_{1}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix},$ (40)
where $L_{1}$ is partitioned as $\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}$ (implicitly, so $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ and $\mathbf{y}_{\text{ini}}(k)=L_{y}\eta_{1}(k)$ from the top block-row of the matrix equation if zero matrix has appropriate dimensions). And we partition $S$ and $L_{f_{1}}$ matrices via:
$S=\begin{bmatrix}S_{u}\\ S_{y}\end{bmatrix},\quad L_{f_{1}}=\begin{bmatrix}L_{f_{u}}\\ L_{f_{y}}\end{bmatrix}$ (41)
3.5 Optimal Multi-Step Predictor
We now apply the parameterization in (40) to a $N$-step ahead prediction of the future output. Optimality in this context refers to finding the minimum variance unbiased predictor of the future output. Using the lower block of (40), we have the following relation for future trajectories:
$\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}S_{u}&L_{f_{u}}\\ S_{y}&L_{f_{y}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix}$ (42)
where it can be expanded for $\mathbf{u}_{N}(k)$ and $\mathbf{y}_{N}(k)$ as:
$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\eta_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (43)
To derive the future control input $\mathbf{u}_{N}(k)$, we substitute the BLUE estimate $\hat{\eta}_{1}(k)$ from Theorem 3.1 into (43), yielding:
$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\hat{\eta}_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (44)
Assuming $L_{f_{u}}$ is invertible, we solve for $\eta_{2}(k)$ from the first equation in (44):
$\eta_{2}(k)=L_{f_{u}}^{-1}\left(\mathbf{u}_{N}(k)-S_{u}\hat{\eta}_{1}(k)\right).$ (45)
Substituting (45) into the expression for $\mathbf{y}_{N}(k)$ in (44), we obtain the predictor for $\mathbf{y}_{N}(k)$ which we denote $\hat{\mathbf{y}}_{N}(k)$:
$\begin{split}\hat{\mathbf{y}}_{N}(k)&=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\left(\mathbf{u}_{N}(k)-S_{u}\hat{\eta}_{1}(k)\right)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\hat{\eta}_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (46)
Using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}{\mathbf{z}}_{\text{ini},m}(k)$:
$\begin{split}\hat{\mathbf{y}}_{N}(k)&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (47)
Define:
$\mathbf{P}_{1}=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})L_{1_{\Sigma_{\zeta}}}^{\dagger},\quad\mathbf{P}_{2}=L_{f_{y}}L_{f_{u}}^{-1},$ (48)
so that:
$\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k).$ (49)
To ensure the mathematical rigor of the NTDPC problem, we introduce the following assumptions and lemma.
Assumption 3.2.
The matrix $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ is invertible, ensuring that the future input $\mathbf{u}_{N}(k)$ uniquely determines $\eta_{2}(k)$ for a given $\hat{\eta}_{1}(k)$.
Assumption 3.3.
The noise $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian with known positive definite covariance $\mathbf{\Sigma}_{\zeta}$. The SVD separates signal and noise such that $\Sigma_{1}$ in (18) captures the dominant system dynamics.
Assumption 3.4.
The constraint sets $\mathcal{U}\subseteq\mathbb{R}^{n_{u}N}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{y}N}$ are convex and non-empty.
Lemma 3.5.
Under Assumptions 3.2 and 3.3, the predictor $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)$ is an unbiased estimator of the true future output $\mathbf{y}_{N}(k)$ that would result from $\mathbf{u}_{N}(k)$ and the true initial state encoded by $\eta_{1}(k)$, i.e., $\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]=S_{y}\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$. The matrices $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ can be computed offline.
Proof.
From (49), $\hat{\mathbf{y}}_{N}(k)$ is derived using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$. By Theorem 3.1, $\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]=\eta_{1}(k)$. The true output corresponding to $\eta_{1}(k)$ and a chosen $\mathbf{u}_{N}(k)$ (which defines $\eta_{2}(k)$) is given by the second line of (43): $\mathbf{y}_{N}^{\text{true}}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)$. With $\eta_{2}(k)=L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$, this becomes $\mathbf{y}_{N}^{\text{true}}(k)=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)$. Taking the expectation of the predictor (47):
$\begin{split}\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)=\mathbf{y}_{N}^{\text{true}}(k).\end{split}$ (50)
Thus, the predictor is unbiased for the true output that would arise from $\eta_{1}(k)$ and $\mathbf{u}_{N}(k)$. The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ follows from their definitions in (48), as they depend on SVD results and system parameters, not online data (except $\mathbf{z}_{\text{ini},m}(k)$ which is an input to the predictor). ∎
We now formulate the NTDPC problem.
Problem 4 [NTDPC]: Given past input and measured output trajectories, NTDPC solves the following optimization problem at each time step $k$:
$\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\sigma_{y}(k)}$ $\quad\|\mathbf{y}_{N}(k)\|_{Q}^{2}+\|\mathbf{u}_{N}(k)\|_{R}^{2}+\|\sigma_{y}(k)\|_{\Lambda_{y}}^{2}$
s.t. $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)+\sigma_{y}(k),$ (51)
$\mathbf{u}_{N}(k)\in\mathcal{U},\quad\hat{\mathbf{y}}_{N}(k)\in\mathcal{Y},$
where $\mathbf{u}_{N}(k)\in\mathbb{R}^{n_{u}N}$ is the input sequence over horizon $N$, $\mathbf{y}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the predicted output sequence in the cost, $\hat{\mathbf{y}}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the constrained output sequence (related to $\mathbf{y}_{N}(k)$ through the model), $\sigma_{y}(k)\in\mathbb{R}^{n_{y}N}$ is a slack variable, $\mathbf{z}_{\text{ini},m}(k)=\text{col}(\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini},m}(k))\in\mathbb{R}^{(n_{u}+n_{y})T_{\text{ini}}}$ contains past inputs and noisy outputs ($\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$), $\mathbf{P}_{1},\mathbf{P}_{2}$ are SVD-derived predictor matrices, $Q,R,\Lambda_{y}\succ 0$ are weighting matrices, and $\mathcal{U},\mathcal{Y}$ are convex constraint sets. Algorithm 1 summarizes the main steps of the NTDPC approach.
The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ reduces online complexity, and the absence of regularization parameters simplifies tuning compared to DeePC, $\gamma$-DDPC, and GDPC.
Remark 3.1.
When calculating the SVD factorization for matrices $\mathbf{Z}_{p}$ and $\mathbf{Z}_{f}V_{2}$ in (12) and (38) in noisy case, it is recommended to normalize the rows of these matrices first, to avoid giving higher priority to input or output. For this purpose, the input and output signals need to be scaled at the first step as $\mathbf{u}(k)\leftarrow M_{u}^{-1}\mathbf{u}(k),\mathbf{y}(k)\leftarrow M_{y}^{-1}\mathbf{y}(k)$ where $M_{u}=\text{diag}(m_{u_{1}},\cdots,m_{u_{n_{u}}})$ and $M_{y}=\text{diag}(m_{y_{1}},\cdots,m_{y_{n_{y}}})$ in which $m_{u_{i}}$ and $m_{y_{i}}$ shows the range of $\mathbf{u}_{i}$ and $\mathbf{y}_{i}$, respectively. In this study, $m_{u_{i}}=\text{mean}(|\mathbf{u}_{i}|);\quad m_{y_{i}}=\text{mean}(|\mathbf{y}_{i}|)$. At the end, the calculated control input $\mathbf{u}(k)$ needs to be denormalized before applying to the plant.
Algorithm 1 Implementation of NTDPC
0:  Generate P.E input and measure output
1:  Normalize the input and output signals
2:  Construct $\mathbf{U}_{p},\mathbf{Y}_{p},\mathbf{U}_{f},\mathbf{Y}_{f}$
3:  Compute and extract $L_{1},V_{1},V_{2},L_{f_{1}}$ using SVD
4:  Calculate the matrices $\mathbf{P}_{1},\mathbf{P}_{2}$
5:  Build $\mathbf{z}_{\text{ini},m}(k)$ using the past $T_{\text{ini}}$ data samples
6:  Obtain $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}~{}{\mathbf{z}}_{\text{ini},m}(k)+\mathbf{P}_{2}~{}\mathbf{u}_{N}(k)$
7:  Solve the QP problem (Problem 4) and obtain $\mathbf{u}_{N}^{*}(k)$
8:  Denormalize the signal $\mathbf{u}_{N}^{*}(k)$
9:  Apply $u(k)=\mathbf{u}_{N_{[0:n_{u}-1]}}^{*}(k)$ to the system
4 Stability Analysis
This section establishes a rigorous stability analysis of the developed NTDPC approach. For both noise-free and noisy cases, we establish stability to the desired equilibrium. In the noise-free case, asymptotic stability ensures tracking of the reference. In the noisy case, input-to-state stability (ISS) guarantees robustness to bounded measurement noise. To facilitate the stability analysis, we define key concepts and notations as done in [24] for DeePC.
Definition 4.1 (Stage Cost).
Assume $Q,R$ are block-diagonal:
$\|\mathbf{y}_{N}(k)\|_{Q}^{2}$ $=\sum_{i=0}^{N-1}\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2},$
$\|\mathbf{u}_{N}(k)\|_{R}^{2}$ $=\sum_{i=0}^{N-1}\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2},$ (52)
where $Q_{i},R_{i}\succ 0$, and $r_{y}\in\mathbb{R}^{n_{y}}$, $r_{u}\in\mathbb{R}^{n_{u}}$ are the steady-state output and input, while $y^{*}(i|k),u^{*}(i|k)$ are the predicted output and input at step $i$ from time $k$. The stage cost is:
$l(y^{*}(i|k),u^{*}(i|k))$ $=\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2}.$
Thus, $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}$.
Definition 4.2 (Non-Minimal State).
The non-minimal state is:
$X_{\text{ini}}(k)$ $=\mathbf{z}_{\text{ini}}(k)=\text{col}(\mathbf{y}_{\text{ini}}(k),\mathbf{u}_{\text{ini}}(k))=\text{col}(y(k-T_{\text{ini}}),\dots,y(k-1),u(k-T_{\text{ini}}),\dots,u(k-1)),$
with $\mathbf{y}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}},\mathbf{u}_{\text{ini}}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}}$. The error state is:
$\tilde{X}_{\text{ini}}(k)$ $=\text{col}(\mathbf{y}_{\text{ini}}(k)-\mathbf{r}_{y},\mathbf{u}_{\text{ini}}(k)-\mathbf{r}_{u}),$
where $\mathbf{r}_{y}=\text{col}(r_{y},\dots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\dots,r_{u})$.
Definition 4.3 (Class $\mathcal{K}$ and $\mathcal{KL}$ Functions).
A function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\varphi(0)=0$. It is $\mathcal{K}_{\infty}$ if $\lim_{s\to\infty}\varphi(s)=\infty$. A function $\beta:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is $\mathcal{KL}$ if $\beta(\cdot,t)\in\mathcal{K}$ for fixed $t\geq 0$, and $\beta(s,t)\to 0$ as $t\to\infty$ for fixed $s>0$. The identity is $\text{id}(s)=s$.
Definition 4.4 (Filtration).
The filtration $\mathcal{F}_{k}$ is the sigma-algebra generated by the measurements $\{\mathbf{z}_{\text{ini},m}(t)\}_{t=0}^{k}$, representing the information available up to time $k$.
4.1 Noise-Free Case
We analyze the stability of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$), tracking references $(r_{y},r_{u})$.
Assumption 4.1 (Positive Definiteness).
There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that for all $(y,u)\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$, where $\mathcal{Y}_{i}\subseteq\mathbb{R}^{n_{y}}$ and $\mathcal{U}_{i}\subseteq\mathbb{R}^{n_{u}}$ are the projections of $\mathcal{Y}$ and $\mathcal{U}$ onto the $i$-th prediction step:
$\alpha_{1}(\|\text{col}(y-r_{y},u-r_{u})\|)\leq l(y,u)\leq\alpha_{2}(\|\text{col}(y-r_{y},u-r_{u})\|),$
where $l(y,u)=\|y-r_{y}\|_{Q_{i}}^{2}+\|u-r_{u}\|_{R_{i}}^{2}$, and $(r_{y},r_{u})\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$ is a constant admissible equilibrium.
Assumption 4.2 (Terminal Stabilizing Condition).
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<\text{id}$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
$l(\hat{y}(N|k),\hat{u}(N|k))\leq(\text{id}-\rho)\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$
Assumption 4.3 (Feasibility).
Problem 4 is feasible for all $k\geq T_{\text{ini}}$ with $\mathbf{\zeta}_{\text{ini}}(k)=0$, and the optimal slack satisfies $\sigma_{y}^{*}(k)=0$. The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
Theorem 4.4.
Under Assumptions 4.1, 4.2, and 4.3, with $\mathbf{\zeta}_{\text{ini}}(k)=0$ and a constant admissible equilibrium $(r_{y},r_{u})$, the closed-loop system under NTDPC is asymptotically stable, i.e., $y(k)\to r_{y}$, $u(k)\to r_{u}$ as $k\to\infty$.
Proof.
Define the non-minimal state as in Definition 4.2:
$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)).$
The error state is $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$. Define the auxiliary function as $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ where the stage cost is $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ with $Q_{i},R_{i}\succ 0$. The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k)),$ since $\sigma_{y}^{*}(k)=0$ in the noise-free case (Assumption 4.3). Consider the storage function [24]:
$V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k)).$
To show positive definiteness of $V(\mathbf{z}_{\text{ini}}(k))$, using Assumption 4.1, we have:
$l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|).$ Since $\tilde{X}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}})-r_{y},\ldots,u(k-1)-r_{u})$, there exists $\alpha_{1,W}\in\mathcal{K}_{\infty}$ such that $W(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|).$ Since $J^{*}(k)\geq 0$, there exists $\alpha_{1,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ For the upper bound:
$l(y(k-i),u(k-i))\leq\alpha_{2}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
so $W(\mathbf{z}_{\text{ini}}(k))\leq T_{\text{ini}}\alpha_{2}\left(\max_{i=1,\dots,T_{\text{ini}}}\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|\right).$ For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)$ (Lemma 3.5) depends on $\mathbf{z}_{\text{ini}}(k)=\tilde{X}_{\text{ini}}(k)+\text{col}(\mathbf{r}_{y},\mathbf{r}_{u})$. Bounded sets $\mathcal{Y},~{}\mathcal{U}$ (Assumption 4.3) and the predictor’s linearity ensure that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$. Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that $J^{*}(k)\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|).$ Combining, there exists $\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Hence, $\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Next, by defining the supply function as:
$s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))=l(\hat{y}(N|k),\hat{u}(N|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})),$
where $\hat{y}(N|k)$, $\hat{u}(N|k)$ satisfy Assumption 4.2. Construct a suboptimal solution at $k+1$:
$\mathbf{u}_{s}(k+1)$ $=\text{col}(u^{*}(1|k),\ldots,u^{*}(N-1|k),\hat{u}(N|k))$
$\mathbf{y}_{s}(k+1)$ $=\text{col}(y^{*}(1|k),\ldots,y^{*}(N-1|k),\hat{y}(N|k)),$ (53)
with $\sigma_{y,s}(k+1)=0$. The suboptimal cost is as $J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k)).$ Since $J^{*}(k)=\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))$ and $J^{*}(k+1)\leq J_{s}(k+1)$:
$J^{*}(k+1)\leq J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))$ (54)
For the auxiliary function $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (55)
Using (61) and (55), thus:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))=[J^{*}(k+1)-J^{*}(k)]+[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))]$ (56)
Then, we can rewrite the preceding equality in (56) as:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))$ $\leq[J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))]$
$\quad-J^{*}(k)+[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))]$
$=s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (57)
By Assumption 4.2, $s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\leq-\rho\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$ Since $l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\geq\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)$, we can rewrite the inequality (57) as:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))\leq-\rho\circ\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)\leq-\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$
Since $V$ is positive definite and radially unbounded with respect to $\tilde{X}_{\text{ini}}(k)$, and satisfies a strict dissipation inequality, standard Lyapunov results [25] imply $\tilde{X}_{\text{ini}}(k)\to 0$ as $k\to\infty$. Thus, $y(k)\to r_{y}$, $u(k)\to r_{u}$, proving asymptotic stability. Since $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ and $W(\mathbf{z}_{\text{ini}}(k))\geq 0$:
$J^{*}(k)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|),$
ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state. This completes the proof. ∎
Remark 4.1.
The bound $J^{*}(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ satisfies the requirement for a $\mathcal{K}_{\infty}$ cost bound, derived within the storage function analysis.
4.2 Noisy Case: Robustness to Measurement Noise
We analyze the stability of NTDPC under measurement noise ($\mathbf{\zeta}_{\text{ini}}(k)\neq 0$), tracking constant admissible references $(r_{y},r_{u})$. The system is shown to be input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$, following the framework in [2, 25].
Assumption 4.5 (Bounded Noise Variance).
There exists $\bar{c}>0$ such that $\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq\bar{c}$ for all $k$, where $\mathcal{F}_{k}$ is the filtration up to time $k$.
Assumption 4.6 (Lipschitz Predictor).
There exists $L_{P}>0$ such that:
$\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k)-\mathbf{z}_{\text{ini}}(k))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k)\|,$
where $L_{P}=\|\mathbf{P}_{1}\|_{2}$ is the spectral norm of $\mathbf{P}_{1}$, and $\mathbf{z}_{\text{ini},m}(k)=\mathbf{z}_{\text{ini}}(k)+\text{col}(0,\mathbf{\zeta}_{\text{ini}}(k))$.
Assumption 4.7 (Terminal Stabilizing Condition - Noisy).
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<id$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(\text{id}-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$
Assumption 4.8 (Feasibility - Noisy).
Problem 4 is feasible with probability at least $1-\epsilon$, for small $\epsilon>0$. There exists $c_{\sigma}>0$ such that:
$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}].$
The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
Remark 4.2.
Assumption 4.7 is satisfied by selecting $\hat{u}(N|k)=r_{u}$ and leveraging the predictor’s robustness (Lemma 3.5). Assumption 4.8 is supported by the sensitivity index $I_{s}\leq 0.7$ (Fig. 5), ensuring noise separability and feasibility.
Theorem 4.9.
Under Assumptions 4.1, 4.5, 4.6, 4.7, and 4.8, with a constant admissible equilibrium $(r_{y},r_{u})$, NTDPC is input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$. There exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
Proof.
Define the non-minimal state as in Definition 4.2:
$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)),$
and the error state $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$. The auxiliary function is $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ with stage cost $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ and $Q_{i},R_{i}\succ 0$. The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}.$ Define the stochastic Lyapunov function:
$V(k)=\mathbb{E}[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$
In order to study positive definiteness of $V(k)$, under Assumption 4.1, we have $l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$ so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|)\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|),$ for some $\alpha_{1,W}\in\mathcal{K}_{\infty}$. For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)+\sigma_{y}^{*}(k)$ and bounded sets $\mathcal{Y},\mathcal{U}$ (Assumption 4.8) imply that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$ and noise effects. By Assumption 4.8:
$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}.$
Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
Combining, there exist $\alpha_{1,V},\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that:
$\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
As dissipation inequality, we construct a suboptimal solution at $k+1$ like (53) with $\sigma_{y,s}(k+1)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k+1)+\mathbf{P}_{2}\mathbf{u}_{s}(k+1)-\mathbf{y}_{s}(k+1)$, where $\hat{u}(N|k),\hat{y}(N|k)$ satisfy Assumption 4.7. The suboptimal cost is:
$J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}.$
By computing $\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]$ as follows:
$\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (58)
Since:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (59)
Then, by subtracting (58) and (59), we have:
$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]$ $=-\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]-\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
$\quad+\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]+\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
By Assumption 4.7:
$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(id-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (60)
For the slack terms, by Assumption 4.6:
$\|\sigma_{y,s}(k+1)\|\leq\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k+1)-\mathbf{z}_{\text{ini}}(k+1))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k+1)\|$ (61)
so, by taking the expectation from both side of (61):
$\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k+1)\|^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\bar{c}.$
Similarly, $\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}$. Let $c_{\sigma,\text{max}}=\max(c_{\sigma},\lambda_{\max}(\Lambda_{y})L_{P}^{2})$. Then:
$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$ (62)
And for $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (63)
By taking the expectation from both side of (63), thus:
$\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]=\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (64)
By combining (62) and (64):
$V(k+1)-V(k)\leq\mathbb{E}[J_{s}(k+1)-J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$ (65)
By substituting (60) in (65), yields:
$V(k+1)-V(k)\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
By Assumption 4.1, there exists $\alpha_{4}\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]\geq\alpha_{4}(\mathbb{E}[\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|])$ (66)
Since $\tilde{X}_{\text{ini}}(k)$ includes these terms and using (66), there exists $\alpha_{4,V}\in\mathcal{K}_{\infty}$ such that:
$V(k+1)-V(k)\leq-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
The term $\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]$ is bounded by the predictor’s output and noise effects, contributing to the ISS gain.
To show ISS property, By Assumption 4.5, $\bar{c}\leq c_{\zeta}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2}$, the dissipation inequality:
$V(k+1)\leq V(k)-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\alpha_{5}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+c_{\zeta,\text{max}}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2},$
where $\alpha_{5}\in\mathcal{K}_{\infty}$ bounds the control cost, and $c_{\zeta,\text{max}}=2c_{\sigma,\text{max}}c_{\zeta}$. Applying stochastic ISS results [25], there exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
Since $V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}$ and $\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]\geq 0$:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c},$
ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state plus a noise-dependent term. Hence, the proof is completed. ∎
Remark 4.3.
The ISS property ensures bounded tracking errors under noise, as validated in Fig. 2, where NTDPC maintains stable tracking of $(r_{y},r_{u})$ with $\sigma_{\zeta}^{2}=0.25I$. The sensitivity index $I_{s}$ (Fig. 5) supports Assumption 4.8, aiding horizon selection.
5 Numerical Simulations
Consider the discrete-time linear model system we aim to control, characterized by the following matrices [26], Section 6.4:
$\begin{split}&A=\begin{bmatrix}0.9997&0.0038&-0.0001&-0.0322\\ -0.0056&0.9648&0.7446&0.0001\\ 0.0020&-0.0097&0.9543&-0.0000\\ 0.0001&-0.0005&0.0978&1.0000\end{bmatrix},\quad\\ &B=\begin{bmatrix}0.0010&0.1000\\ -0.0615&0.0183\\ -0.1133&0.0586\\ -0.0057&0.0029\end{bmatrix},C=\begin{bmatrix}1.0000&0&0&0\\ 0&-1.0000&0&7.7400\end{bmatrix}\end{split}$ (67)
These matrices are obtained by discretizing the dynamics of the longitudinal motion of a Boeing 747 using a zero-order hold with a sampling time $T_{s}=0.1\text{ s}$. The system has two inputs: the throttle $u_{1}$ and the elevator angle $u_{2}$, and two outputs: the aircraft’s longitudinal velocity and climb rate.
The control objective is to increase the longitudinal velocity ($y_{1}(t)$) from 0 ft/s to 10 ft/s while maintaining the climb rate ($y_{2}(t)$) at zero. This corresponds to a reference output $r_{y}=\begin{bmatrix}10&0\end{bmatrix}^{T}$. The inputs and outputs are constrained as follows:
$\begin{split}&U:=\left\{u\in\mathbb{R}^{2}:\begin{bmatrix}-20\\ -20\end{bmatrix}\leq u\leq\begin{bmatrix}20\\ 20\end{bmatrix}\right\},\quad\\ &Y:=\left\{y\in\mathbb{R}^{2}:\begin{bmatrix}-25\\ -15\end{bmatrix}\leq y\leq\begin{bmatrix}25\\ 15\end{bmatrix}\right\}.\end{split}$ (68)
In this simulation, we set $T_{\text{ini}}=N=20$, data length for Hankel matrices $M=2500$.
5.1 Noise-free case
The Fig.1 illustrates the performance of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$) when tracking constant admissible references ($(r_{y},r_{u})$), as analyzed in Section 5.1. The results are consistent with the asymptotic stability proven in Theorem 4.4, where $y(k)\to r_{y}$ and $u(k)\to r_{u}$ as $k\to\infty$ under Assumptions 4.1, 4.2, and 4.3. The top plot displays the primary output $y_{1}(k)$ for NTDPC (blue), SPC (red), and SMMPC (green) over time, with the dashed line representing the reference $r_{y}$; all methods converge to $r_{y}$, but NTDPC exhibits the fastest transient response and minimal overshoot, reflecting the effectiveness of the data-driven predictor and terminal condition. The second plot shows a secondary output $y_{2}(k)$, with similar convergence behavior, where NTDPC (blue) stabilizes more quickly than SPC (red) and SMMPC (green), supporting the robustness of the NTDPC framework to initial conditions encoded in $\mathbf{z}_{\text{ini}}(k)$. The third plot presents the primary input $u_{1}(k)$, where NTDPC (blue) adjusts rapidly from an initial value, stabilizing near $r_{u}$ (dashed line) with less oscillation compared to SPC (red) and SMMPC (green), consistent with the bounded input constraints in $\mathcal{U}$. Finally, the bottom plot depicts a secondary input $u_{2}(k)$, showing NTDPC (blue) achieving a steady state closer to $r_{u}$ with reduced variability, further validating the stability properties derived from the Lyapunov function $V(\mathbf{z}_{\text{ini}}(k))$. Overall, the convergence of outputs and inputs to their respective references demonstrates the asymptotic stability of NTDPC, as predicted by the storage function analysis $V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))$, and the superior performance of NTDPC over SPC and SMMPC highlights the advantage of the noise-tolerant predictor and optimization framework, particularly in the absence of noise.
Figure 1: Output trajectories for noise-free case
5.2 Noisy case
In Fig. 2, the simulation results demonstrate the performance of NTDPC in comparison to SPC and SMMPC under the noise effect, $\sigma_{\zeta}^{2}=0.25I$. While SMMPC exhibits aggressive and unstable control actions leading to poor regulation of outputs, NTDPC and SPC show reasonable results, despite an initial overshoot and undershoot, ultimately achieves stable tracking of the reference for both outputs with damped oscillations and less aggressive control inputs that settle over time. This balanced response in both outputs and the settling behavior of the control inputs highlight the effectiveness and stability of the NTDPC and SPC strategies compared to the oscillatory and poorly regulated responses observed with SMMPC.
Figure 2: Output trajectories over 50 Monte Carlo runs for noisy case
To quantify controller performance, the objective function is evaluated and plotted in Fig. 3 which reveals the superiority of NTDPC and SPC compared to SMMPC. It is worthy to note that there is a slight difference between NTDPC and SPC throughout the entire simulation, highlighting the effectiveness of the designed NTDPC. In summary, the figures provide a comprehensive comparison of NTDPC, SPC and SMMPC control strategies, highlighting their performance in terms of output and input trajectories and their respective variability. This comparison aids in understanding the strengths and weaknesses of each method in maintaining desired control objectives amidst system noise and constraints.
Figure 3: Performance index over 50 Monte Carlo runs for noisy case
5.3 Sensitivity Evaluation
Increasing horizon lengths improves estimation quality but increases computational cost and adaptation time in time-varying systems. We propose a sensitivity index to guide parameter selection under various noise conditions. When using noisy data, matrix $\Sigma_{2}$ in (18) becomes non-zero. Ideally, $\Sigma_{2}$’s singular values should be noise-related and significantly smaller than those of $\Sigma_{1}$. In practice, both matrices contain signal and noise components. Figure 4 shows singular value patterns for $\Sigma_{1}$ and $\Sigma_{2}$ with $\sigma_{\zeta}^{2}=0.01I$ and $T_{\text{ini}}=N=15$, with a vertical line separating the two sets.
Figure 4: Singular value distribution for $T_{\text{ini}}=N=15$ and $\sigma_{\zeta}^{2}=0.01I$
To quantify signal-noise separability, we propose:
$I_{s}=\frac{\sigma_{\text{max}}^{2}(\Sigma_{2})}{\sigma_{\text{min}}^{2}(\Sigma_{1})}$ (69)
Lower $I_{s}$ values indicate better signal-noise separability. Figure 5 shows $I_{s}$ behavior as $T_{\text{ini}}$ increases from 10 to 50 with noise variances from 0.01 to 0.32. Empirically, algorithm convergence requires $I_{s}\lesssim 0.7$. This confirms higher horizons are needed for higher noise variances and guides horizon selection. Each point represents the mean from 50 Monte Carlo simulations, with shading indicating one standard deviation. Since these analyses can be performed before controller implementation (or offline for LTI systems), this approach provides practical guidance for horizon selection.
Figure 5: Sensitivity index for different noises and $T_{\text{ini}}$ values
5.4 Computational Complexity
In this section, we will evaluate computational complexity of the developed NTDPC compared to SPC. In the context of a time-varying system, the prediction model needs to be updated at each sample time. Comparing the computational effort for updating this model between SPC and the NTDPC, focusing solely on the more costly calculations of pseudoinverse and SVD, reveals the following for a simplified SISO system with input/output horizons $T_{\text{ini}}=N=T_{h}$ (using $T_{h}$ to avoid confusion with data length $T$) and a data window length $M$:
For SPC, estimating the prediction matrices involves computing the pseudoinverse of the Hankel matrix $\begin{bmatrix}\mathbf{U}_{p}&\mathbf{Y}_{p}&\mathbf{U}_{f}\end{bmatrix}^{T}$ (or its non-transposed version depending on formulation), which has dimensions $3T_{h}\times M$. Assuming $M\gg 3T_{h}$, the computational complexity of the pseudoinverse is approximately $\mathcal{O}((3T_{h})^{2}M)=\mathcal{O}(9T_{h}^{2}M)$.
For the NTDPC method, the matrix estimation per sample time primarily involves two SVD computations: one for the matrix $Z_{p}$ (dimensions $2T_{h}\times M$) and another for the matrix $Z_{f}V_{2}$ (dimensions approximately $2T_{h}\times(M-T_{h}-n)$, where $n$ is the system order, here simplified to $2T_{h}\times(M-T_{h})$). The computational complexity of the SVDs are approximately $\mathcal{O}((2T_{h})^{2}M)$ and $\mathcal{O}((2T_{h})^{2}(M-T_{h}-n))$, respectively, assuming the number of rows is less than the number of columns in both cases ($2T_{h}<M$ and $2T_{h}<M-T_{h}-n$). Summing these, the total dominant SVD cost for NTDPC is approximately $\mathcal{O}(4T_{h}^{2}M+4T_{h}^{2}(M-T_{h}-n))\approx\mathcal{O}(8T_{h}^{2}M)$ for large $M$.
This comparison of the dominant factorization/inversion costs highlights that NTDPC offers a reduction in computational effort per sample time with a lower constant factor (approximately 8 compared to 9) in the $\mathcal{O}(T_{h}^{2}M)$ complexity. Furthermore, similar to other hybrid methods the NTDPC works with reduced order Hankel matrices which in turn reduces the memory needed for the online matrix estimation step, contributing to improved overall efficiency for time-varying systems.
6 Conclusion
This paper presents a comprehensive analysis of the Noise-Tolerant Data-Driven Predictive Control (NTDPC) framework, addressing the critical challenge of measurement noise in hybrid data-driven predictive control while achieving robust tracking of non-zero references. By leveraging singular value decomposition (SVD), NTDPC effectively separates system dynamics from noise within reduced-order Hankel matrices, enabling accurate trajectory predictions with shorter data horizons and reduced computational complexity. The stability analysis demonstrates that NTDPC achieves asymptotic stability in the noise-free case, ensuring precise tracking of the reference equilibrium. In the presence of measurement noise, NTDPC exhibits input-to-state stability (ISS), guaranteeing robust performance with bounded tracking errors proportional to the noise magnitude.
Numerical simulations on a flight control benchmark validate the theoretical findings, showcasing NTDPC’s superior performance compared to SMMPC and SPC. The output trajectories demonstrate stable tracking with damped oscillations and less aggressive control inputs, while the performance index highlights NTDPC’s effectiveness in minimizing tracking errors. The proposed sensitivity index, $I_{s}$, provides a practical tool for selecting prediction horizons under varying noise conditions, with empirical results indicating that $I_{s}\lesssim 0.7$ ensures algorithm convergence. This index, computed offline for linear time-invariant systems, enhances the applicability of NTDPC in real-world scenarios. Furthermore, the computational complexity analysis reveals that NTDPC reduces the dominant factorization costs by approximately 11% compared to SPC, with lower memory requirements due to reduced-order Hankel matrices, making it well-suited for time-varying systems.

### v2

@@
 DPC approaches can be divided into two categories: indirect and direct methods [5, 6, 7].
-Indirect methods, exemplified by Subspace Predictive Control (SPC) [8], first identify a prediction model from data, which is subsequently used as an constraint in the optimal control problem.
+Indirect methods, exemplified by Subspace Predictive Control (SPC) [8], first identify a prediction model from data, which is subsequently used as a constraint in the optimal control problem.
 This intermediate modeling phase makes these approaches particularly suitable for integration with large datasets.
@@
 In contrast, direct DPC approaches, such as Data-Enabled Predictive Control (DeePC) [10], instead of estimating a prediction model incorporate the input output samples directly in the optimal control problem [11].
-This is possible by introducing a regularization vector "$\mathbf{g}(k)$" based on Willems’ fundamental lemma [12].
+This is possible by introducing a regularization vector $\mathbf{g}(k)$" based on Willems’ fundamental lemma [12].
 This approach enables the predictor to be "control-oriented" specifically focused on generating predictions for control purposes, however, the optimization problem has a larger size as $\mathbf{g}(k)$ is part of the decision variables that needs to be optimized [13].
@@
 This is even more important when knowing that SPC as a base of comparison is efficient for noisy cases as well.
-To improve the noise handling of hybrid approaches, this study proposes a different for estimation of $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ that takes the noise characteristics into consideration.
+To improve the noise handling of hybrid approaches, this study proposes a different approach for estimation of $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ that takes into account the noise characteristics.
 Compared to the mentioned hybrid methods, our approach borrows the assumption of orthogonality of components from [17] and replaces the LQ factorization with SVD which can exctract the system dynamics from the noisy data when estimating the $\mathbf{g}(k)$ components.
-The idea of using SVD in data driven predictive control is not new as [18, 11, 19] have used this decomposition.
+The application of Singular Value Decomposition (SVD) in data-driven control has recently gained attention, primarily for computational model order reduction within direct control formulations [18, 11, 19].
+However, the potential of SVD to geometrically separate signal and noise subspaces within a hybrid predictive framework remains underexplored.
+This study uses SVD beyond simple dimensionality reduction, utilizing it as a robust filtering mechanism.
+By embedding this projection into the predictor construction, the proposed NTDPC framework systematically replaces the heuristic regularization typical of DeePC with a rigorous, noise-tolerant estimation step.
 In [18], using SVD the length of past data and hence the dimension of Hankel matrix is reduced.
@@
 2 PRELIMINARIES
-Consider a discrete-time linear dynamical system influenced by zero-mean Gaussian noise $\zeta(k)\sim N(0,\sigma^{2}_{\zeta}I)$:
+Consider a discrete-time linear dynamical system influenced by zero-mean Gaussian noise $\zeta(k)\sim\mathcal{N}(0,\Sigma_{\zeta})$:
 $x(k+1)$ $=Ax(k)+Bu(k),\quad k\in\mathbb{N},$ (1)
@@
 We assume that $(A,B)$ is controllable and $(A,C)$ is observable.
-By applying a persistently exciting input sequence $\{u(k)\}_{k\in\mathbb{N}[0,T]}$ of length $T$ to system (1), we obtain a corresponding output sequence $\{y(k)\}_{k\in\mathbb{N}[0,T]}$.
-For an input-output model, we introduce the past horizon $T_{\text{ini}}$ that limits the window of past input-output data necessary to compute the current output:
-$y(k)=\sum_{i=1}^{T_{\text{ini}}}a_{i}y(k-i)+\sum_{i=1}^{T_{\text{ini}}}b_{i}u(k-i),$ (3)
-Consider a system whose input and output data can be described, for some real-valued matrix coefficients $a_{i}$ and $b_{i}$.
-For simplicity of notation, we assume a uniform past data window length, denoted by $T_{\text{ini}}$, for both the input and output signals.
-Now, for any starting time instant $k\geq 0$ within the available data vector and for any data vector length $j\geq 1$, we define:
+By applying a persistently exciting input sequence $\{u(k)\}_{k=0}^{T-1}$ of length $T$ to system (1), we obtain a corresponding output sequence $\{y(k)\}_{k=0}^{T-1}$.
+The measurement noise $\zeta(k)$ is assumed to be i.i.d., zero-mean Gaussian, and uncorrelated in time.
+For an input-output model, we introduce the past horizon $T_{\text{ini}}$ to characterize the system’s memory.
+Let $y_{o}(k)$ denote the underlying noise-free output of the system (i.e., the output of (1)-(2) if $\zeta(k)\equiv 0$).
+The system dynamics and the measurement model are described by:
+$y_{o}(k)$ $=\sum_{i=1}^{T_{\text{ini}}}a_{i}y_{o}(k-i)+\sum_{i=1}^{T_{\text{ini}}}b_{i}u(k-i),$ (3a)
+$y(k)$ $=y_{o}(k)+\zeta(k).$ (3b)
+In this representation, the coefficients $a_{i}$ and $b_{i}$ define the system dynamics.
+For a general MIMO system, these are real-valued matrices with dimensions $a_{i}\in\mathbb{R}^{n_{y}\times n_{y}}$ and $b_{i}\in\mathbb{R}^{n_{y}\times n_{u}}$, while for a SISO system, they reduce to scalar parameters.
+Furthermore, for simplicity of notation, we assume the same memory horizon $T_{\text{ini}}$ for both the input and output regressors, rather than assigning distinct lag lengths to $u$ and $y$.
+Now, for any starting time instant $k\geq 0$ within the available data vector and for any data vector length $j\geq 1$, we define the trajectory vectors:
 $\bar{u}(k,j)$ $:=\operatorname{col}(u(k),\dots,u(k+j-1)),$
-$\bar{y}(k,j)$ $:=\operatorname{col}(y(k),\dots,y(k+j-1)).$
-Lets define $N$ as the future horizon and using $T_{\text{ini}}$, we construct the following sequences of past and future inputs and outputs:
-$\begin{split}&\mathbf{u}_{\text{ini}}(k):=\text{col}(u(k-T_{\text{ini}}),\ldots,u(k-1)),\quad\\ &\mathbf{y}_{\text{ini}}(k):=\text{col}(y(k-T_{\text{ini}}+1),\ldots,y(k)),\\ &\mathbf{u}_{N}(k):=\text{col}(u(k),\ldots,u(k+N)),\\ &\mathbf{y}_{N}(k):=\text{col}(y(k),\ldots,y(k+N)),\end{split}$ (4)
-To transition from the model-based prediction to data-driven counterpart, we assume access to an informative set of input-output pairs $\{\bar{u},\bar{y}\}$ of length $T$.
-Informativity is ensured by the persistent excitation of the input sequence, which must be of order $T_{\text{ini}}+N+n$, implies that $T$ must be sufficiently long, specifically $T\geq n_{u}(T_{\text{ini}}+N+n).$ Let $N\geq T_{\text{ini}}$, the Hankel matrices can be defined as follows:
-$\mathbf{U}_{p}$ $,\displaystyle:=[\bar{u}(0,T_{\text{ini}}),$$\dots,\bar{u}(T-1,T_{\text{ini}})],$ (5)
-$\mathbf{Y}_{p}$ $,\displaystyle:=[\bar{y}(1,T_{\text{ini}}),$$\dots,\bar{y}(T,T_{\text{ini}})],$
-$\mathbf{U}_{f}$ $,\displaystyle:=[\bar{u}(T_{\text{ini}},N),$$\dots,\bar{u}(T_{\text{ini}}+T-1,N)],$
-$\mathbf{Y}_{f}$ $,\displaystyle:=[\bar{y}(T_{\text{ini}}+1,N),$$\dots,\bar{y}(T_{\text{ini}}+T,N)].$
-which form the foundation of the data-driven predictive control approach.
-In the next, we introduce three different DPC strategies, by directly using input–output data instead of an explicit model of the system and a state estimator.
-Problem 1 [SPC]: Subspace Predictive Control (SPC), represents an early indirect data-driven MPC technique.
-It directly incorporates results from subspace identification into the MPC design by estimating prediction matrices from input/output data based on the relationship $\mathbf{y}(k)=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)&\mathbf{y}_{\text{ini}}(k)&\mathbf{u}(k)\end{bmatrix}^{T}.$ The unknown matrix $\mathbf{\Theta}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}$ is estimated by solving the least-squares problem $\min_{\mathbf{\Theta}}\left\|\mathbf{Y}_{f}-\mathbf{\Theta}\begin{bmatrix}\mathbf{U}_{p}&\mathbf{Y}_{p}&\mathbf{U}_{f}\end{bmatrix}^{T}\right\|_{\text{Fro}}^{2}$, with solution $\mathbf{\Theta}^{*}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\hat{\Gamma}}\end{bmatrix}$.
-Then, the SPC can be formulated as [8]:
+$\bar{y}(k,j)$ $:=\operatorname{col}(y(k),\dots,y(k+j-1)).$ (4)
+In this notation, the first argument $k$ denotes the starting time instant, and the second argument $j$ explicitly denotes the length of the data window.
+Let us define $N$ as the future prediction horizon. Using $T_{\text{ini}}$, we construct the following sequences of past and future inputs and outputs:
+$\begin{split}\mathbf{u}_{\text{ini}}(k)&:=\text{col}(u(k-T_{\text{ini}}),\ldots,u(k-1)),\\ \mathbf{y}_{\text{ini}}(k)&:=\text{col}(y(k-T_{\text{ini}}+1),\ldots,y(k)),\\ \mathbf{u}_{N}(k)&:=\text{col}(u(k),\ldots,u(k+N-1)),\\ \mathbf{y}_{N}(k)&:=\text{col}(y(k),\ldots,y(k+N-1)).\end{split}$ (5)
+To transition from model-based prediction to a data-driven counterpart, we assume access to an informative set of input-output pairs $\{\bar{u},\bar{y}\}$ of length $T$.
+Informativity is ensured by the persistent excitation of the input sequence, where the rank of $\mathbf{Z}_{p}$ matrix defined in (11) must be higher than $T_{\text{ini}}+N+n$.
+This implies that the data length $T$ must be sufficiently long, specifically $T\geq(n_{u}+1)(T_{\text{ini}}+N+n)-1$.
+The persistency of excitation condition is imposed on the input sequence $\{\bar{u}\}$ , which guarantees that the corresponding Hankel matrix $\mathbf{Z}_{p}$ constructed from the input-output data has sufficient rank.
+Let $M=T-T_{ini}-N+1$ be the number of columns in the Hankel matrices. The matrices are defined as:
+$\mathbf{U}_{p}$ $=[\bar{u}(0,T_{ini}),\dots,\bar{u}(M-1,T_{ini})],$ (6)
+$\mathbf{Y}_{p}$ $=[\bar{y}(0,T_{ini}),\dots,\bar{y}(M-1,T_{ini})],$
+$\mathbf{U}_{f}$ $=[\bar{u}(T_{ini},N),\dots,\bar{u}(T_{ini}+M-1,N)],$
+$\mathbf{Y}_{f}$ $=[\bar{y}(T_{ini},N),\dots,\bar{y}(T_{ini}+M-1,N)].$
+These matrices form the foundation of the data-driven predictive control approach.
+Next, we introduce three different DPC strategies that utilize this data directly, avoiding the need for an explicit system model or state estimator.
+Problem 1 [SPC]: Subspace Predictive Control (SPC) represents an early indirect data-driven MPC technique.
+It directly incorporates results from subspace identification into the MPC design by estimating prediction matrices from input/output data based on the relationship $\mathbf{y}_{N}(k)=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)^{T}&\mathbf{y}_{\text{ini}}(k)^{T}&\mathbf{u}_{N}(k)^{T}\end{bmatrix}^{T}$.
+The unknown matrix $\mathbf{\Theta}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}$ is estimated by solving the least-squares problem $\min_{\mathbf{\Theta}}\left\|\mathbf{Y}_{f}-\mathbf{\Theta}\begin{bmatrix}\mathbf{U}_{p}^{T}&\mathbf{Y}_{p}^{T}&\mathbf{U}_{f}^{T}\end{bmatrix}^{T}\right\|_{\text{Fro}}^{2}$, yielding the solution $\mathbf{\Theta}^{*}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\hat{\Gamma}}\end{bmatrix}$.
+The SPC problem is then formulated as [8]:
 $\begin{split}\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}\\ \text{subject to}\quad&\mathbf{y}_{N}(k)=\mathbf{P}_{1}\mathbf{u}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{y}_{\text{ini}}(k)+\mathbf{\hat{\Gamma}}\mathbf{u}_{N}(k),\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}.\end{split}$ (7)
-Unlike SPC, which relies on estimating prediction matrices, Data-Enabled Predictive Control (DeePC) is a purely data-driven control approach that directly predicts the future output sequence $\mathbf{y}(k)$ from Hankel matrix data.
-Next, the DeePC problem is defined as follows:
+Unlike SPC, which relies on an intermediate identification step to find prediction matrices, DeePC is a direct approach that embeds the raw data samples into the optimization problem.
 Problem 2 [DeePC]: Given immediate past input and measured output signals, find a future input sequence by solving [20]:
 $\begin{split}\min_{\Theta(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}+\lambda_{g}l_{g}(\mathbf{g}(k))+\lambda_{y}\|\sigma_{y}(k)\|_{2}^{2}\\ \text{subject to}\quad&\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\mathbf{g}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)+\sigma_{y}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix},\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}\end{split}$ (8)
-where $\Theta(k)=\{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\mathbf{g}(k),\sigma_{y}(k)\}$ and, $l_{g}(\mathbf{g}(k))$ is a user-defined regularization function over $\mathbf{g}(k)$, often chosen as $\|\mathbf{g}(k)\|_{2}^{2}$, leading to the term "DeePC".
-The variable $\sigma_{y}(k)$ is introduced to prevent infeasibility if $\{\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini}}(k)\}$ is not in the behavioral set, especially in the presence of disturbances.
-To address the need for regularization in the predictive control problem, a variant of DPC was recently developed [17], employing an optimal (minimum variance) unbiased predictor for future output trajectories.
-Problem 3 [SMMPC]: Given immediate past input and measured output signals, find a future input sequence by solving [17]:
+where $\Theta(k)=\{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\mathbf{g}(k),\sigma_{y}(k)\}$ and $l_{g}(\mathbf{g}(k))$ is a user-defined regularization function over $\mathbf{g}(k)$, often chosen as $\|\mathbf{g}(k)\|_{2}^{2}$.
+The variable $\sigma_{y}(k)$ is a slack variable introduced to ensure feasibility in the presence of noise.
+While DeePC offers a direct data-driven formulation, the optimization vector $\mathbf{g}(k)$ scales with the dataset size, which can be computationally intensive.
+To mitigate this and improve noise handling, hybrid approaches have been proposed that pre-compute an optimal predictor offline.
+Problem 3 [SMMPC]: A variant of DPC, Signal Matrix Model Predictive Control (SMMPC) [17], employs a minimum-variance unbiased predictor derived offline. The control problem is solved as:
 $\begin{split}\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}\\ \text{subject to}\quad&\mathbf{y}_{N}(k)=\mathbf{E}_{1}\mathbf{u}_{\text{ini}}(k)+\mathbf{E}_{2}\mathbf{y}_{\text{ini}}(k)+\mathbf{E}_{3}\mathbf{u}_{N}(k),\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}.\end{split}$ (9)
-The faster execution is attributed to the offline computation of $\mathbf{E}_{1}$, $\mathbf{E}_{2}$, and $\mathbf{E}_{3}$.
+The computational efficiency of SMMPC arises from the offline calculation of the predictor matrices $\mathbf{E}_{1}$, $\mathbf{E}_{2}$, and $\mathbf{E}_{3}$, which replaces the high-dimensional online optimization of $\mathbf{g}(k)$ required in DeePC.
 3 Indirect DPC using reduced-order Hankel matrices
-Data-driven MPC with Hankel matrices often has more columns than rows, leading to non-unique $\mathbf{g}(k)$, requiring regularization of its norm.
-DeePC’s large matrices can delay dynamics capture in time-varying systems.
-Smith et.al, [17] separates $\mathbf{g}(k)$ identification into two phases for a smaller control problem: $\mathbf{g}_{\text{ini}}(k)$ from past data (using LQ factorization for uniqueness due to non-full rank Hankel matrix) and $\mathbf{g}_{\text{f}}(k)$ from future data in the null space of the past Hankel matrix.
-Inspired by [17], we decompose $\mathbf{g}(k)$ into orthogonal parts:
-$\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\big{)}=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}.$ (9)
-While sharing the idea of separate $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ identification using past and future data, we address SMMPC’s noise handling limitation.
-Noise can make Hankel matrices full row rank, causing LQ factorization to prioritize noise over signal, significantly impacting SMMPC performance with noisy data.
-This study focuses on introducing constraints for solution uniqueness and minimizing sensor noise impact.
-We propose SVD as a superior factorization method for handling noisy data, detailed in the following section.
-SMMPC uses LQ factorization to uniquely determine $\mathbf{g}_{\text{ini}}(k)$ from partitioned $\mathbf{U}_{p}$ and $\mathbf{Y}_{p}$, and $\mathbf{g}_{f}(k)$ in their null space using future data.
-While creative for deterministic cases, this may be suboptimal with noisy data.
-We propose a systematic method to handle noise in both the output signals and the Hankel matrices.
-By leveraging singular value decomposition (SVD), we decompose the Hankel matrices into components associated with system dynamics and noise, respectively, based on the assumption that the noise variance is smaller than the signal variance.
-This decomposition enables a robust estimation of the system dynamics.
-A key contribution of our approach is the replacement of the high-dimensional regularization vector $\mathbf{g}(k)=\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)$ with lower-dimensional vectors $\eta_{1}(k)$ and $\eta_{2}(k)$, exploiting the rank deficiency of the Hankel matrices.
-Unlike [17], which uses LQ factorization to separately solve deterministic equations, our method formulates the estimation of $\eta_{1}(k)$ as a constrained optimization problem that simultaneously accounts for deterministic input constraints and stochastic output measurements.
-This joint optimization is necessary because the deterministic and stochastic components are coupled in our SVD-based formulation.
-The resulting NTDPC problem is formulated as a convex quadratic program (QP), which is computationally efficient due to offline computation of key matrices and eliminates the need for user-defined regularization parameters, unlike DeePC [20], $\gamma$-DDPC [21, 22], and GDPC [23].
+In data-driven predictive control, the Hankel matrices constructed from historical data typically possess significantly more columns than rows ($M\gg(n_{u}+n_{y})T_{ini}$).
+This structural characteristic renders the linear system for the regularization vector $\mathbf{g}(k)$ underdetermined, admitting an infinite set of possible solutions.
+Consequently, standard approaches require regularization terms (e.g., minimum norm) to select a specific solution.
+To address the underdetermined nature of the problem, we adopt a sequential identification strategy that decomposes the regularization vector $\mathbf{g}(k)$ into two orthogonal components: a past component $\mathbf{g}_{\text{ini}}(k)$ determined by historical data, and a future component $\mathbf{g}_{\text{f}}(k)$ that governs the predictive behavior.
+This decomposition is formulated as:
+$\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\big(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\big)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}$.
+(10)
+Equation (10) poses an identification challenge due to the rank properties of the data matrix.
+In the noise-free case, the Hankel matrix is rank-deficient (rank equal to the system order plus the input horizon), reflecting the underlying low-dimensional dynamics.
+However, measurement noise invariably renders the matrix full row rank.
+In this full-rank scenario, standard approaches like LQ factorization force a unique solution that effectively fits the noise, treating stochastic perturbations as valid system dynamics.
+To prevent this overfitting, we propose a transformation that enforces the underlying low-rank structure.
+By assuming a spectral gap, we utilize SVD to separate the dominant signal subspace ($\Sigma_{1}$) from the noise subspace ($\Sigma_{2}$).
+This allows us to project the high-dimensional $\mathbf{g}(k)$ onto a lower-dimensional latent basis ($\eta_{1},\eta_{2}$), converting the ill-posed noisy problem into a well-posed, noise-tolerant estimation problem.
 3.1 Derivation of a Parsimonious Formulation
-We condense the mapping in (9) into a minimal parameterization. Consider stacked Hankel matrices representing the past and future input-output sequences:
-$\mathbf{Z}_{p}:=\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\end{bmatrix},\quad\mathbf{Z}_{f}:=\begin{bmatrix}\mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix},\quad\mathbf{z}_{\text{ini}}(k):=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\end{bmatrix}$ (10)
-Using SVD, we factorize $\mathbf{Z}_{p}$ as:
-$\mathbf{Z}_{p}=W\Sigma V^{T}$ (11)
-where $W$ and $V$ are unitary matrices. If the system is noise free (i.e. $\zeta=0$), $\Sigma$ is not full rank and we have
-$\mathbf{Z}_{p}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}=L_{1}V_{1}^{T}$ (12)
-in which $L_{1}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}\\ 0\end{bmatrix}$, and $\Sigma_{1}$ is a diagonal matrix containing non-zero singular entries.
-According to [17], the rank of $\mathbf{Z}_{p}$ is equal to $n_{u}T_{\text{ini}}+n$ and therefore $\Sigma_{1}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times(n_{u}T_{\text{ini}}+n)}$ and $V_{1}^{T}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times M_{c}}$, where $M_{c}$ is the number of columns in $\mathbf{Z}_{p}$.
-(Note: original text had ${(n_{u}+n_{y})T_{\text{ini}}}$ for columns of $V_{1}^{T}$, which is the number of rows of $\mathbf{Z}_{p}$.
-$V_{1}^{T}$ has $M_{c}$ columns, the same as $\mathbf{Z}_{p}$.
-Its rows are $n_{u}T_{\text{ini}}+n$.
-Please verify this dimension.)
-Now, from (9) we have:
-$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\Big{)}\end{split}$ (13)
-where assuming that $\mathbf{g}_{f}(k)$ is in null space of $\mathbf{Z}_{p}$, one can write
-$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)\end{split}$ (14)
-By substituting $\mathbf{Z}_{p}$ from (12) into (14), yields
-$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}V_{1}^{T}\mathbf{g}_{\text{ini}}(k)\end{split}$ (15)
-Defining a vector $\eta_{1}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}+n}$ as $\eta_{1}(k)=V_{1}^{T}\mathbf{g}_{\text{ini}}(k)$, then we will obtain the following least-squares problem as:
-$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k),\end{split}$ (16)
-where $\eta_{1}(k)$ will be later calculated by solving an estimation problem.
-Since $L_{1}$ has full column rank (as $\Sigma_{1}$ is invertible), we can recover $\mathbf{g}_{\text{ini}}(k)$ from $\eta_{1}(k)$ as:
-$\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ (17)
-In case $\zeta\neq 0$, i.e. $\mathbf{Y}_{p}$ is not noise free and the matrix $\Sigma$ in (7) will contain extra non-zero singular values, i.e.:
-$\Sigma=\begin{bmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{bmatrix}$ (18)
-However, assuming that the variance of noise is smaller than the variation of inputs and outputs, the SV decomposition by nature would separate the noise from the signal.
-Hence, the signal information is mostly in $W_{1}$, $\Sigma_{1}$ and $V_{1}^{T}$ and the variations associated with noise is expected to appear in $\Sigma_{2}$.
+We condense the mapping in (10) into a minimal parameterization. Consider the stacked Hankel matrices representing the past and future input-output sequences:
+$\mathbf{Z}_{p}:=\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\end{bmatrix},\quad\mathbf{Z}_{f}:=\begin{bmatrix}\mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix},\quad\mathbf{z}_{\text{ini}}(k):=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\end{bmatrix}$.
+(11)
+$Z_{p}$Past Data==
+$W$
+[…]
+$\Sigma_{1}$
+$\sigma_{1}$$\sigma_{r}$$\Sigma_{2}\ll$00SignalNoise
+$V^{\top}$
+[…]
+[…]
+Figure 1: SVD Decomposition of the noisy Hankel matrix $Z_{p}$.
+$\Sigma_{1}$ captures dominant dynamics; $\Sigma_{2}$ represents noise.
+$W$ is partitioned into signal columns $W_{1}$ and noise columns $W_{2}$.
+By performing SVD factorization on the past data matrix $\mathbf{Z}_{p}$:
+$\mathbf{Z}_{p}=W\Sigma V^{T}.$ (12)
+where $W$ and $V$ are unitary matrices. In the ideal noise-free case ($\zeta=0$), the matrix $\Sigma$ is rank-deficient. We can partition the decomposition as:
+$\mathbf{Z}_{p}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}=W_{1}\Sigma_{1}V_{1}^{T}=L_{1}V_{1}^{T},$ (13)
+where we define $L_{1}:=W_{1}\Sigma_{1}$.
+Here, $\Sigma_{1}$ is a diagonal matrix containing the non-zero singular values (see Fig.1).
+The rank of the noiseless $\mathbf{Z}_{p}$ is given by $r=n_{u}T_{\text{ini}}+n$ [21].
+Consequently, the dimensions are defined as $\Sigma_{1}\in\mathbb{R}^{r\times r}$ and $V_{1}^{T}\in\mathbb{R}^{r\times M}$, where $M$ is the number of columns in $\mathbf{Z}_{p}$.
+In cases of low Signal-to-Noise Ratio (SNR), the singular values of the system dynamics and noise can indeed overlap or "flip," making simple truncation problematic.
+Therefore, we explicitly clarify that our method operates under the assumption of a sufficient spectral gap between the dominant system dynamics and the noise floor.
+Substituting this decomposition into the past trajectory equation (10):
+$\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\big(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\big).$ (14)
+We enforce the condition that $\mathbf{g}_{\text{f}}(k)$ lies in the null space of $\mathbf{Z}_{p}$, such that $\mathbf{Z}_{p}\mathbf{g}_{\text{f}}(k)=0$. This simplifies (14) to:
+$\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)=L_{1}V_{1}^{T}\mathbf{g}_{\text{ini}}(k).$ (15)
+We now define the latent vector $\eta_{1}(k)\in\mathbb{R}^{r}$ as the projection of the regularization vector onto the active subspace:
+$\eta_{1}(k)=V_{1}^{T}\mathbf{g}_{\text{ini}}(k).$ (16)
+This yields the following linear system describing the initial trajectory:
+$\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k).$ (17)
+Since we seek the minimum-norm solution for $\mathbf{g}_{\text{ini}}(k)$ that satisfies the data consistency, we restrict $\mathbf{g}_{\text{ini}}(k)$ to the row space of $\mathbf{Z}_{p}$ (spanned by $V_{1}$).
+This allows us to recover $\mathbf{g}_{\text{ini}}(k)$ uniquely from $\eta_{1}(k)$ via:
+$\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k).$ (18)
+In the realistic case where $\zeta\neq 0$, the measured output matrix $\mathbf{Y}_{p}$ is noisy, and $\mathbf{Z}_{p}$ becomes full rank.
+Let $M$ be the number of columns.
+By utilizing SVD, yielding a rectangular singular value matrix $\Sigma\in\mathbb{R}^{(n_{u}+n_{y})T_{ini}\times M}$, which is partitioned as:
+$\Sigma=\begin{bmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{bmatrix}$, (19)
+where $\Sigma_{1}\in\mathbb{R}^{r\times r}$ contains the $r$ dominant singular values associated with the system dynamics.
+The second block, $\Sigma_{2}\in\mathbb{R}^{((n_{u}+n_{y})T_{ini}-r)\times(M-r)}$, is a rectangular matrix containing the remaining singular values dominated by noise (padded with zeros as necessary).
+By truncating $\Sigma_{2}$, we obtain a low-rank approximation that serves as a robust basis for estimating $\eta_{1}(k)$.
 3.2 Best Linear Unbiased Estimator for $\eta_{1}(k)$
-We now formalize the estimation of $\eta_{1}(k)$ in the presence of noise using the BLUE framework.
+We now formalize the estimation of the latent vector $\eta_{1}(k)$ in the presence of noise.
+This formulation distinguishes between the deterministic inputs (controller outputs) and the stochastic measurements (noisy sensor measurements) using the Best Linear Unbiased Estimator (black) framework.
+In the following derivation, we treat the matrix $L_{y}$ as a deterministic basis for the system dynamics.
+This is justified by the offline SVD truncation step (Section 3.1), which uses the large data history $M$ to filter out measurement noise and extract the dominant signal subspace.
+Consequently, while the offline raw data is noisy, the resulting basis $L_{y}$ represents the denoised nominal model, whereas the single online measurement $z_{ini,m}(k)$ retains significant noise that must be filtered.
 Theorem 3.1.
-Let the measured initial condition be $\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)^{\top}&\mathbf{y}_{\text{ini},m}(k)^{\top}\end{bmatrix}^{\top}$, where $\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, and $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian noise with covariance $\mathbf{\Sigma}_{\zeta}$.
-Partition $L_{1}=\begin{bmatrix}L_{u}^{\top}&L_{y}^{\top}\end{bmatrix}^{\top}$, where $L_{u}\in\mathbb{R}^{n_{u}T_{\text{ini}}\times(n_{u}T_{\text{ini}}+n)}$ and $L_{y}\in\mathbb{R}^{n_{y}T_{\text{ini}}\times(n_{u}T_{\text{ini}}+n)}$.
-The best linear unbiased estimator (BLUE) of $\eta_{1}(k)$ is given by:
-$\begin{split}\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)\end{split}$ (19)
-where:
-$L_{1_{\Sigma_{\zeta}}}^{\dagger}=\begin{bmatrix}L_{u}^{\dagger}&\left(I-L_{u}^{\dagger}L_{u}\right)\left(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\right)^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\end{bmatrix}.$ (20)
-and $L_{u}^{\dagger}=L_{u}^{T}(L_{u}L_{u}^{T})^{-1}$ is the pseudo-inverse of $L_{u}$.
+Let the measured initial trajectory be $\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)^{\top}&\mathbf{y}_{\text{ini},m}(k)^{\top}\end{bmatrix}^{\top}$, where $\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\boldsymbol{\zeta}_{\text{ini}}(k)$, with measurement noise $\boldsymbol{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$, and $\mathbf{z}_{\text{ini},m}(k)=L_{1}\eta_{1}(k)$.
+By partitioning the system basis matrix $L_{1}$ from (13) as $L_{1}=\begin{bmatrix}L_{u}^{\top}&L_{y}^{\top}\end{bmatrix}^{\top}$, where $L_{u}\in\mathbb{R}^{n_{u}T_{\text{ini}}\times r}$ and $L_{y}\in\mathbb{R}^{n_{y}T_{\text{ini}}\times r}$, the Best Linear Unbiased Estimator (black) of $\eta_{1}(k)$ is given by:
+$\hat{\eta}_{1}(k)=\mathcal{K}\mathbf{z}_{\text{ini},m}(k),$ (20)
+where the estimator gain matrix $\mathcal{K}\in\mathbb{R}^{r\times(n_{u}+n_{y})T_{\text{ini}}}$ is defined as:
+$\mathcal{K}=\begin{bmatrix}L_{u}^{\dagger}-K_{y}L_{y}L_{u}^{\dagger}&K_{y}\end{bmatrix}$, (21)
+with the auxiliary correction gain $K_{y}$ given by:
+$K_{y}=P\left(P^{\top}L_{y}^{\top}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}P\right)^{\dagger}P^{\top}L_{y}^{\top}\mathbf{\Sigma}_{\zeta}^{-1},$ (22)
+and $P=I-L_{u}^{\dagger}L_{u}$ is the orthogonal projector onto the null space of $L_{u}$.
 Proof.
-The measured initial condition is:
-$\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini},m}(k)\end{bmatrix}=\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}\eta_{1}(k)+\begin{bmatrix}0\\ \mathbf{\zeta}_{\text{ini}}(k)\end{bmatrix},$ (21)
-where $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ is deterministic, and $\mathbf{y}_{\text{ini},m}(k)=L_{y}\eta_{1}(k)+\mathbf{\zeta}_{\text{ini}}(k)$ is stochastic with $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$.
-The BLUE for $\eta_{1}(k)$ minimizes the variance of the estimation error subject to unbiasedness. This translates to the constrained least-squares problem:
-$\begin{split}\hat{\eta}_{1}(k)&=\arg\min_{\eta_{1}(k)}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}(k)\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}(k)\right)\\ &\hskip 14.22636pt\text{subject to:}~{}~{}~{}~{}~{}~{}~{}L_{u}\eta_{1}(k)=\mathbf{u}_{\text{ini}}(k).\end{split}$ (22)
-To solve this, we use Lagrange multipliers. Define the Lagrangian (omitting $(k)$ for brevity in derivation):
-$\begin{split}\mathcal{L}(\eta_{1},\lambda)&=\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)+2\lambda^{T}\left(L_{u}\eta_{1}-\mathbf{u}_{\text{ini}}\right).\end{split}$ (23)
-Taking partial derivatives with respect to $\eta_{1}$ and $\lambda$, and setting them to zero, we get:
-$\frac{\partial\mathcal{L}}{\partial\eta_{1}}$ $=-2L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)+2L_{u}^{T}\lambda=0,$ (24)
-$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=L_{u}\eta_{1}-\mathbf{u}_{\text{ini}}=0.$ (25)
-From (24), we obtain:
-$L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\eta_{1}=L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda.$ (26)
-The general solution to $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}$ from (25) is:
-$\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}+\left(I-L_{u}^{\dagger}L_{u}\right)\xi,$ (27)
-where $\xi$ is an arbitrary vector. Substitute (27) into (26):
-$\begin{split}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\left(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}+\left(I-L_{u}^{\dagger}L_{u}\right)\xi\right)=L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda.\end{split}$ (28)
-To find $\lambda$, multiply by $L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}$ (assuming $(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})$ is invertible, which is typical if $L_{y}$ has full column rank and $\mathbf{\Sigma}_{\zeta}$ is positive definite).
-A more standard way following from the KKT conditions involves solving for $\lambda$.
-From (26), if $L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}$ is invertible:
-$\eta_{1}=(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda)$ (29)
-Substitute this into the constraint $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}$:
-$L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda)=\mathbf{u}_{\text{ini}}$ (30)
-Then solve for $\lambda$:
-$\begin{split}\lambda=&[L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{u}^{T}]^{-1}[L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-\mathbf{u}_{\text{ini}}]\end{split}$ (31)
-Substituting $\lambda$ back into (29) gives $\hat{\eta}_{1}$.
-The result (19) with (20) is a known solution for this type of constrained least-squares problem (mixed deterministic-stochastic model).
-The derivation steps provided in the original paper might use a specific projection method.
-For brevity here, we acknowledge the standard result.
-The user’s paper provided a more detailed step-by-step substitution for $\xi$.
-I’ll restore the user’s flow:
-From (25), we have $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}(k)$. The general solution for $\eta_{1}$ satisfying this is:
-$\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi,$ (32)
-where $\xi$ is an arbitrary vector.
-Substitute (32) into the first KKT condition (24) (after rearranging it to solve for $\eta_{1}$ if $L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}$ is invertible, or substitute into objective): The objective is to minimize $J(\eta_{1})=\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)$.
-Substitute $\eta_{1}$ from (32):
-$\begin{split}J(\xi)=&\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi)\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\\ &\times\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi)\right)\end{split}$ (33)
-Let $\tilde{\mathbf{y}}=\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)$ and $\tilde{L}_{y}=L_{y}(I-L_{u}^{\dagger}L_{u})$.
-Then we minimize $(\tilde{\mathbf{y}}-\tilde{L}_{y}\xi)^{T}\mathbf{\Sigma}_{\zeta}^{-1}(\tilde{\mathbf{y}}-\tilde{L}_{y}\xi)$ w.r.t $\xi$.
-The solution is $\hat{\xi}=(\tilde{L}_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{L}_{y})^{\dagger}\tilde{L}_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{\mathbf{y}}$.
-Assuming $(L_{y}(I-L_{u}^{\dagger}L_{u}))^{T}\mathbf{\Sigma}_{\zeta}^{-1}(L_{y}(I-L_{u}^{\dagger}L_{u}))$ is invertible, which simplifies to $((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}$.
-Then $\hat{\xi}=((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}(I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k))$.
-Substituting this $\hat{\xi}$ into (32) yields:
-$\begin{split}\hat{\eta}_{1}(k)=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)&+\left(I-L_{u}^{\dagger}L_{u}\right)((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}\\ &\times(I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)\right).\end{split}$ (34)
-This is equivalent to the solution given in (19) using the definition in (20) by recognizing $(I-L_{u}^{\dagger}L_{u})$ as a projector, and that for a projector $P$, $P(P^{T}AP)^{-1}P^{T}=(A_{p})^{-1}$ where $A_{p}$ is A projected on the space.
-The form in the paper, $(I-L_{u}^{\dagger}L_{u})(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}$, assumes that $(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})$ is invertible and that this formulation correctly finds the $\xi$ that minimizes the cost within the null space of $L_{u}$.
-This structure is standard.
-To express this compactly using the definition in (20): $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$. This estimator is unbiased since:
-$\mathbb{E}[\hat{\eta}_{1}(k)]=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbb{E}[\mathbf{z}_{\text{ini},m}(k)]=L_{1_{\Sigma_{\zeta}}}^{\dagger}\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}\eta_{1}(k).$ (35)
-For this to be $\eta_{1}(k)$, we need $L_{1_{\Sigma_{\zeta}}}^{\dagger}\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}=I$. Using (20):
-$L_{u}^{\dagger}L_{u}+(I-L_{u}^{\dagger}L_{u})(L_{y}^{T}\Sigma_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\Sigma_{\zeta}^{-1}L_{y}=L_{u}^{\dagger}L_{u}+I-L_{u}^{\dagger}L_{u}=I$
-.
-This confirms unbiasedness.
-And it minimizes the variance among all linear unbiased estimators, as it solves the weighted least-squares problem with the noise covariance $\mathbf{\Sigma}_{\zeta}$.
-This completes the proof.
-∎
+The estimation problem is formulated as a constrained generalized least-squares optimization.
+We seek the estimate $\hat{\eta}_{1}(k)$ that minimizes the variance of the measurement residual subject to the deterministic input constraints:
+$\min_{\eta_{1}}$ $J=\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)^{\top}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)$ (23)
+s.t. $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}(k).$
+The general solution satisfying the linear constraint is $\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+P\xi$, where $\xi$ is an arbitrary vector in the null space of $L_{u}$.
+Substituting this parameterization into the objective function $J$ transforms the problem into an unconstrained minimization over $\xi$:
+$\min_{\xi}\left\|\mathbf{\Sigma}_{\zeta}^{-1/2}\left(\tilde{\mathbf{y}}(k)-L_{y}P\xi\right)\right\|_{2}^{2},$ (24)
+where $\tilde{\mathbf{y}}(k)=\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)$ represents the output residual after removing the effect of the known inputs.
+The optimal $\xi$ is obtained via the weighted pseudoinverse:
+$\hat{\xi}=\left(L_{y}P\right)^{\dagger}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{\mathbf{y}}(k)=\left(P^{\top}L_{y}^{\top}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}P\right)^{\dagger}P^{\top}L_{y}^{\top}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{\mathbf{y}}(k).$ (25)
+Substituting $\hat{\xi}$ back into the expression for $\eta_{1}$ yields:
+$\hat{\eta}_{1}(k)=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+P\hat{\xi}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+K_{y}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)\right).$ (26)
+Grouping the terms multiplying $\mathbf{u}_{\text{ini}}(k)$ and $\mathbf{y}_{\text{ini},m}(k)$ results in the block matrix form presented in (21).
+We assume that the output subspace $L_{y}P$ has sufficient rank to identify the free components of $\eta_{1}$.
+Crucially, this estimation relies strictly on the available information set $\mathcal{F}_{k}$ (the past noisy trajectory).
+While the offline Hankel matrix contains future data correlations, the online estimator must remain causal; therefore, $\hat{\eta}_{1}(k)$ is derived solely from $\mathbf{z}_{\text{ini},m}(k)$ without utilizing future block rows, which are unavailable at time $k$.
+Remark 3.1 (Estimation and Robustness of $\Sigma_{\zeta}$).
+While $\Sigma_{\zeta}$ is assumed known, in practice it can be estimated via the sample covariance of steady-state sensor data or through an iterative feasible generalized least squares (FGLS) approach using residuals.
+The proposed estimator $\hat{\eta}_{1}(k)$ exhibits inherent robustness to estimation errors as explained below:
+  1. 1.
+    Scale Invariance: The estimator is invariant to absolute scaling of the covariance matrix (i.e., replacing $\Sigma_{\zeta}$ with $\alpha\Sigma_{\zeta}$ for $\alpha>0$ yields the identical gain $\mathcal{K}$).
+Consequently, for MIMO systems the exact noise magnitude estimation is not required; only the relative variance ratios between sensor channels are necessary.
+  2. 2.
+    Unbiasedness: An incorrect covariance structure results in a loss of minimum-variance optimality (efficiency) but preserves the unbiasedness of the predictor, as the projection condition holds for any positive definite weighting matrix.
 3.3 Future Trajectory Parameterization
-To find $\mathbf{g}_{f}(k)$, note that all $\mathbf{g}(k)$ (likely meant $\mathbf{g}_{f}(k)$) in the null-space of $\mathbf{Z}_{p}$ with the condition $\mathbf{g}_{f}\perp\mathbf{g}_{\text{ini}}$ are given by
-$\mathbf{g}_{f}(k)=V_{2}x_{np}(k)$ (36)
-where $x_{np}(k)\in\mathbb{R}^{M_{c}-(n_{u}T_{\text{ini}}+n)}$ (Note: $M$ was used, $M_{c}$ is num columns of $\mathbf{Z}_{p}$.
-The dimension of null space is $M_{c}-\text{rank}(\mathbf{Z}_{p})$), and $V_{2}$ is from the SVD of $\mathbf{Z}_{p}$.
-The future dynamics satisfy:
-$\mathbf{Z}_{f}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)\Big{)}=\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}$ (37)
-Now, multiplying $\mathbf{Z}_{f}$ by both sides of (36) (this seems to be a typo, equation should be $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=\mathbf{Z}_{f}V_{2}x_{np}(k)$) and factorizing $\mathbf{Z}_{f}V_{2}$ using SVD as follows:
-$\begin{split}\mathbf{Z}_{f}V_{2}&=W_{f}\Sigma_{f}V_{f}^{T}=\begin{bmatrix}W_{f_{1}}&W_{f_{2}}\end{bmatrix}\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{f_{1}}^{T}\\ V_{f_{2}}^{T}\end{bmatrix}=W_{f_{1}}\Sigma_{f_{1}}V_{f_{1}}^{T}\end{split}$ (38)
-where $L_{f_{1}}=W_{f_{1}}\Sigma_{f_{1}}$. Define $\eta_{2}(k)=V_{f_{1}}^{T}x_{np}(k)\in\mathbb{R}^{n_{u}N}$, so:
-$\mathbf{Z}_{f}\mathbf{g}_{f}(k)=\mathbf{Z}_{f}V_{2}x_{np}(k)=L_{f_{1}}\eta_{2}(k).$ (39)
-In the noisy case, $\Sigma_{f}=\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&\Sigma_{f_{2}}\end{bmatrix}$, with noise variations primarily in $\Sigma_{f_{2}}$.
+To determine the future component of the regularization vector, denoted as $\mathbf{g}_{f}(k)$, any valid solution must lie in the null-space of $\mathbf{Z}_{p}$ to satisfy the orthogonality condition $\mathbf{g}_{f}\perp\mathbf{g}_{\text{ini}}$.
+Consequently, $\mathbf{g}_{f}(k)$ can be parameterized as:
+$\mathbf{g}_{f}(k)=V_{2}x_{np}(k),$ (27)
+where $V_{2}$ is the matrix of right singular vectors corresponding to the null space of $\mathbf{Z}_{p}$, derived from the SVD in (13).
+The vector $x_{np}(k)\in\mathbb{R}^{M-r}$ represents the degrees of freedom in the null space.
+The future input-output dynamics satisfy the fundamental data equation:
+$\mathbf{Z}_{f}\Big(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)\Big)=\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}$. (28)
+Substituting (27) into the future term $\mathbf{Z}_{f}\mathbf{g}_{f}(k)$ yields $\mathbf{Z}_{f}V_{2}x_{np}(k)$.
+To compress this representation, we perform an SVD on the matrix product $\mathbf{Z}_{f}V_{2}$:
+$\mathbf{Z}_{f}V_{2}=W_{f}\Sigma_{f}V_{f}^{T}=\begin{bmatrix}W_{f_{1}}&W_{f_{2}}\end{bmatrix}\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{f_{1}}^{T}\\ V_{f_{2}}^{T}\end{bmatrix}=W_{f_{1}}\Sigma_{f_{1}}V_{f_{1}}^{T}.$ (29)
+Since $\mathbf{Z}_{f}V_{2}$ represents the free future evolution consistent with past data, its dominant singular directions capture the controllable future dynamics.
+Let $L_{f_{1}}=W_{f_{1}}\Sigma_{f_{1}}$.
+We define a reduced coordinate vector $\eta_{2}(k)\in\mathbb{R}^{n_{u}N}$ as $\eta_{2}(k)=V_{f_{1}}^{T}x_{np}(k)$, where under sufficient excitation of future inputs, the rank of $\mathbf{Z}_{f}V_{2}$ equals $n_{u}N$.
+This allows us to express the contribution of the future regularization vector compactly as:
+$\mathbf{Z}_{f}\mathbf{g}_{f}(k)=L_{f_{1}}\eta_{2}(k).$ (30)
+In the presence of noise, the matrix $\Sigma_{f}$ will be full rank, taking the form $\Sigma_{f}=\text{diag}(\Sigma_{f_{1}},\Sigma_{f_{2}})$.
+However, consistent with our noise-handling strategy, we assume the dominant singular values in $\Sigma_{f_{1}}$ capture the system dynamics, while $\Sigma_{f_{2}}$ is dominated by noise and is therefore truncated.
+Note that $\eta_{2}(k)\in\mathbb{R}^{n_{u}N}$ and $L_{f_{1}}\in\mathbb{R}^{(n_{u}+n_{y})N\times n_{u}N}$.
+The matrix $L_{f_{1}}$ is partitioned as $L_{f_{1}}=\begin{bmatrix}L_{f_{u}}^{\top}&L_{f_{y}}^{\top}\end{bmatrix}^{\top}$ with $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ and $L_{f_{y}}\in\mathbb{R}^{n_{y}N\times n_{u}N}$.
 3.4 Dimension Reduction of the Future Trajectory Parameterization
-To summarize the equations, $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ can be related to $\eta_{1}(k)$ and $\eta_{2}(k)$ using (17) and (39) (via $x_{np}(k)$).
-Substituting these into (9): The past part is $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$.
-Using $\mathbf{Z}_{p}\mathbf{g}_{f}(k)=0$ and (16), then $\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k)$.
-The future part is $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$.
-Using $\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ from (17) and $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=L_{f_{1}}\eta_{2}(k)$ from (39).
-So, $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}V_{1}\eta_{1}(k)+L_{f_{1}}\eta_{2}(k)$.
-By defining $S=\mathbf{Z}_{f}V_{1}$, the past and future input/output data can be written in terms of intermediate variables $\eta_{1}(k)$ and $\eta_{2}(k)$ as follows:
-$\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ S&L_{f_{1}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix},$ (40)
-where $L_{1}$ is partitioned as $\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}$ (implicitly, so $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ and $\mathbf{y}_{\text{ini}}(k)=L_{y}\eta_{1}(k)$ from the top block-row of the matrix equation if zero matrix has appropriate dimensions).
-And we partition $S$ and $L_{f_{1}}$ matrices via:
-$S=\begin{bmatrix}S_{u}\\ S_{y}\end{bmatrix},\quad L_{f_{1}}=\begin{bmatrix}L_{f_{u}}\\ L_{f_{y}}\end{bmatrix}$ (41)
+We now unify the past and future parameterizations into a single low-order system.
+The past trajectory is given by $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$.
+Since $\mathbf{g}_{f}(k)$ lies in the null space of $\mathbf{Z}_{p}$, this simplifies to $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)$, which we previously derived in (17) as $\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k)$.
+For the future trajectory, we substitute $\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ and the result from (30) into (28):
+$\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}V_{1}\eta_{1}(k)+L_{f_{1}}\eta_{2}(k).$ (31)
+Defining the coupling matrix $S=\mathbf{Z}_{f}V_{1}$, we can express the full trajectory vector in terms of the latent variables $\eta_{1}(k)$ and $\eta_{2}(k)$:
+$\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ S&L_{f_{1}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix}$.
+(32)
+To facilitate the predictor derivation, we partition the matrices $L_{1}$, $S$, and $L_{f_{1}}$ to separate input and output components:
+$L_{1}=\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix},\quad S=\begin{bmatrix}S_{u}\\ S_{y}\end{bmatrix},\quad L_{f_{1}}=\begin{bmatrix}L_{f_{u}}\\ L_{f_{y}}\end{bmatrix}$. (33)
+where the dimensions of the partitioned matrices are defined consistent with the signal dimensions: $L_{u}\in\mathbb{R}^{n_{u}T_{ini}\times r}$ and $L_{y}\in\mathbb{R}^{n_{y}T_{ini}\times r}$, $S_{u}\in\mathbb{R}^{n_{u}N\times r}$, $S_{y}\in\mathbb{R}^{n_{y}N\times r}$.
+This implicitly satisfies $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ and $\mathbf{y}_{\text{ini}}(k)=L_{y}\eta_{1}(k)$.
 3.5 Optimal Multi-Step Predictor
-We now apply the parameterization in (40) to a $N$-step ahead prediction of the future output.
-Optimality in this context refers to finding the minimum variance unbiased predictor of the future output.
-Using the lower block of (40), we have the following relation for future trajectories:
-$\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}S_{u}&L_{f_{u}}\\ S_{y}&L_{f_{y}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix}$ (42)
-where it can be expanded for $\mathbf{u}_{N}(k)$ and $\mathbf{y}_{N}(k)$ as:
-$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\eta_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (43)
-To derive the future control input $\mathbf{u}_{N}(k)$, we substitute the BLUE estimate $\hat{\eta}_{1}(k)$ from Theorem 3.1 into (43), yielding:
-$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\hat{\eta}_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (44)
-Assuming $L_{f_{u}}$ is invertible, we solve for $\eta_{2}(k)$ from the first equation in (44):
+We apply the parameterization in (32) to construct an $N$-step ahead predictor.
+Our goal is to derive the minimum variance unbiased predictor of the future output $\mathbf{y}_{N}(k)$ given the estimated initial state and future control inputs.
+Expanding the lower block of (32):
+$\begin{split}\mathbf{u}_{N}(k)&=S_{u}\eta_{1}(k)+L_{f_{u}}\eta_{2}(k),\\ \mathbf{y}_{N}(k)&=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k).\end{split}$ (34)
+Substituting the black estimate $\hat{\eta}_{1}(k)$ from Theorem 3.1 into (34) yields the prediction model:
+$\begin{split}\mathbf{u}_{N}(k)&=S_{u}\hat{\eta}_{1}(k)+L_{f_{u}}\eta_{2}(k),\\ \mathbf{y}_{N}(k)&=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}\eta_{2}(k).\end{split}$ (35)
+Since $L_{f_{u}}$ is invertible, we can uniquely solve for the future latent vector $\eta_{2}(k)$ using the specified future inputs $\mathbf{u}_{N}(k)$:
 $\eta_{2}(k)=L_{f_{u}}^{-1}\left(\mathbf{u}_{N}(k)-S_{u}\hat{\eta}_{1}(k)\right).$ (36)
-Substituting (45) into the expression for $\mathbf{y}_{N}(k)$ in (44), we obtain the predictor for $\mathbf{y}_{N}(k)$ which we denote $\hat{\mathbf{y}}_{N}(k)$:
+Substituting (36) into the output equation in (35), we obtain the explicit predictor $\hat{\mathbf{y}}_{N}(k)$:
 $\begin{split}\hat{\mathbf{y}}_{N}(k)&=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\left(\mathbf{u}_{N}(k)-S_{u}\hat{\eta}_{1}(k)\right)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\hat{\eta}_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (37)
-Using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}{\mathbf{z}}_{\text{ini},m}(k)$:
-$\begin{split}\hat{\mathbf{y}}_{N}(k)&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (47)
-Define:
-$\mathbf{P}_{1}=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})L_{1_{\Sigma_{\zeta}}}^{\dagger},\quad\mathbf{P}_{2}=L_{f_{y}}L_{f_{u}}^{-1},$ (48)
-so that:
+Replacing $\hat{\eta}_{1}(k)$ with the estimator $\mathcal{K}~\mathbf{z}_{\text{ini},m}(k)$ derived in the Theorem 3.1:
+$\hat{\mathbf{y}}_{N}(k)=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathcal{K}\mathbf{z}_{\text{ini},m}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).$ (38)
+We define the predictor system matrices as:
+$\mathbf{P}_{1}=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})\mathcal{K},\quad\mathbf{P}_{2}=L_{f_{y}}L_{f_{u}}^{-1},$ (39)
+where $P_{1}\in\mathbb{R}^{n_{y}N\times(n_{u}+n_{y})T_{ini}}$, $P_{2}\in\mathbb{R}^{n_{y}N\times n_{u}N}$. Then, it leads to the compact affine prediction model:
 $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k).$ (40)
-To ensure the mathematical rigor of the NTDPC problem, we introduce the following assumptions and lemma.
+To ensure the mathematical rigor of the NTDPC formulation, we introduce the following assumptions and lemma.
 Assumption 3.2.
-The matrix $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ is invertible, ensuring that the future input $\mathbf{u}_{N}(k)$ uniquely determines $\eta_{2}(k)$ for a given $\hat{\eta}_{1}(k)$.
+[Future Excitation and Reduced-Order Parameterization] The future input sequence $u_{N}$ is sufficiently exciting such that
+$\operatorname{rank}(Z_{f}V_{2})=n_{u}N,$
+where $V_{2}$ spans the nullspace of $Z_{p}$.
+As a result, the reduced future latent variable $\eta_{2}\in\mathbb{R}^{n_{u}N}$ provides a complete parameterization of all admissible future trajectories consistent with the past data.
+Remark 3.2.
+The reduced future input mapping $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ is assumed to be full rank in the noise-free case, which is guaranteed provided the offline training data satisfies the condition of persistency of excitation.
+In the presence of noise or numerical degeneracy, a regularized pseudoinverse of $L_{f_{u}}$ is employed to compute the future latent variable $\eta_{2}$.
+This guarantees the existence and uniqueness of the multi-step predictor in practice.
 Assumption 3.3.
-The noise $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian with known positive definite covariance $\mathbf{\Sigma}_{\zeta}$.
-The SVD separates signal and noise such that $\Sigma_{1}$ in (18) captures the dominant system dynamics.
-Assumption 3.4.
-The constraint sets $\mathcal{U}\subseteq\mathbb{R}^{n_{u}N}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{y}N}$ are convex and non-empty.
+The constraint sets $\mathcal{U}\subseteq\mathbb{R}^{n_{u}N}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{y}N}$ are convex, closed, and non-empty.
 Lemma 3.4.
-Under Assumptions 3.2 and 3.3, the predictor $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)$ is an unbiased estimator of the true future output $\mathbf{y}_{N}(k)$ that would result from $\mathbf{u}_{N}(k)$ and the true initial state encoded by $\eta_{1}(k)$, i.e., $\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]=S_{y}\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$.
-The matrices $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ can be computed offline.
+Under Assumptions 3.2 and invertible property of $L_{f_{u}}$, the predictor in (40) is an unbiased estimator of the true future output $\mathbf{y}_{N}^{\text{true}}(k)$ given the parameters $\eta_{1}(k)$ and $\mathbf{u}_{N}(k)$.
+That is, $\mathbb{E}[\hat{\mathbf{y}}_{N}(k)\mid\eta_{1}(k),\mathbf{u}_{N}(k)]=\mathbf{y}_{N}^{\text{true}}(k)$.
 Proof.
-From (49), $\hat{\mathbf{y}}_{N}(k)$ is derived using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$.
-By Theorem 3.1, $\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]=\eta_{1}(k)$.
-The true output corresponding to $\eta_{1}(k)$ and a chosen $\mathbf{u}_{N}(k)$ (which defines $\eta_{2}(k)$) is given by the second line of (43): $\mathbf{y}_{N}^{\text{true}}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)$.
-With $\eta_{2}(k)=L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$, this becomes $\mathbf{y}_{N}^{\text{true}}(k)=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)$.
-Taking the expectation of the predictor (47):
-$\begin{split}\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)=\mathbf{y}_{N}^{\text{true}}(k).\end{split}$ (50)
-Thus, the predictor is unbiased for the true output that would arise from $\eta_{1}(k)$ and $\mathbf{u}_{N}(k)$.
-The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ follows from their definitions in (48), as they depend on SVD results and system parameters, not online data (except $\mathbf{z}_{\text{ini},m}(k)$ which is an input to the predictor).
-∎
-We now formulate the NTDPC problem.
-Problem 4 [NTDPC]: Given past input and measured output trajectories, NTDPC solves the following optimization problem at each time step $k$:
+The true output of the system, given the latent state $\eta_{1}(k)$ and input $\mathbf{u}_{N}(k)$, is defined by the second line of (34): $\mathbf{y}_{N}^{\text{true}}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)$.
+Substituting the exact solution for $\eta_{2}(k)$, we get:
+$\mathbf{y}_{N}^{\text{true}}(k)=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).$ (41)
+Taking the expectation of the predictor (38), and noting that $\mathbf{u}_{N}(k)$ is deterministic and $\mathbb{E}[\hat{\eta}_{1}(k)\mid\eta_{1}(k)]=\eta_{1}(k)$ (from Theorem 3.1):
+$\begin{split}\mathbb{E}[\hat{\mathbf{y}}_{N}(k)]&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathbb{E}[\hat{\eta}_{1}(k)]+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\mathbf{y}_{N}^{\text{true}}(k).\end{split}$ (42)
+This confirms the predictor is unbiased. Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ rely only on the SVD of the offline data, they can be computed offline. ∎
+We now formally state the NTDPC optimization problem.
+Problem 4 [NTDPC]: Given past input and measured output trajectories, NTDPC computes the optimal control sequence by solving the following convex optimization problem at each time step $k$:
 $\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\sigma_{y}(k)}\quad$ $\|\mathbf{y}_{N}(k)\|_{Q}^{2}+\|\mathbf{u}_{N}(k)\|_{R}^{2}+\|\sigma_{y}(k)\|_{\Lambda_{y}}^{2}$
-s.t. $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)+\sigma_{y}(k),$ (51)
-$\mathbf{u}_{N}(k)\in\mathcal{U},\quad\hat{\mathbf{y}}_{N}(k)\in\mathcal{Y},$
-where $\mathbf{u}_{N}(k)\in\mathbb{R}^{n_{u}N}$ is the input sequence over horizon $N$, $\mathbf{y}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the predicted output sequence in the cost, $\hat{\mathbf{y}}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the constrained output sequence (related to $\mathbf{y}_{N}(k)$ through the model), $\sigma_{y}(k)\in\mathbb{R}^{n_{y}N}$ is a slack variable, $\mathbf{z}_{\text{ini},m}(k)=\text{col}(\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini},m}(k))\in\mathbb{R}^{(n_{u}+n_{y})T_{\text{ini}}}$ contains past inputs and noisy outputs ($\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$), $\mathbf{P}_{1},\mathbf{P}_{2}$ are SVD-derived predictor matrices, $Q,R,\Lambda_{y}\succ 0$ are weighting matrices, and $\mathcal{U},\mathcal{Y}$ are convex constraint sets.
-Algorithm 1 summarizes the main steps of the NTDPC approach.
-The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ reduces online complexity, and the absence of regularization parameters simplifies tuning compared to DeePC, $\gamma$-DDPC, and GDPC.
+subject to $\mathbf{y}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)+\sigma_{y}(k),$ (43)
+$\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y},$
+where $\mathbf{y}_{N}(k)\in\mathbb{R}^{n_{y}N}$ represents the predicted future output used in the cost function, and $\sigma_{y}(k)\in\mathbb{R}^{n_{y}N}$ is a slack variable ensuring feasibility against noise and model mismatch.
+The weighting matrices $Q,R,\Lambda_{y}\succ 0$ are user-defined.
+The vector $\mathbf{z}_{\text{ini},m}(k)$ contains the most recent $T_{\text{ini}}$ input and noisy output measurements.
+Algorithm 1 summarizes the implementation steps.
 Remark 3.3.
-When calculating the SVD factorization for matrices $\mathbf{Z}_{p}$ and $\mathbf{Z}_{f}V_{2}$ in (12) and (38) in noisy case, it is recommended to normalize the rows of these matrices first, to avoid giving higher priority to input or output.
-For this purpose, the input and output signals need to be scaled at the first step as $\mathbf{u}(k)\leftarrow M_{u}^{-1}\mathbf{u}(k),\mathbf{y}(k)\leftarrow M_{y}^{-1}\mathbf{y}(k)$ where $M_{u}=\text{diag}(m_{u_{1}},\cdots,m_{u_{n_{u}}})$ and $M_{y}=\text{diag}(m_{y_{1}},\cdots,m_{y_{n_{y}}})$ in which $m_{u_{i}}$ and $m_{y_{i}}$ shows the range of $\mathbf{u}_{i}$ and $\mathbf{y}_{i}$, respectively.
-In this study, $m_{u_{i}}=\text{mean}(|\mathbf{u}_{i}|);\quad m_{y_{i}}=\text{mean}(|\mathbf{y}_{i}|)$.
-At the end, the calculated control input $\mathbf{u}(k)$ needs to be denormalized before applying to the plant.
+To prevent numerical bias during the SVD factorization of $\mathbf{Z}_{p}$ and $\mathbf{Z}_{f}V_{2}$, it is recommended to normalize the input and output data.
+This balances the singular values when inputs and outputs have significantly different magnitudes.
+We apply the scaling:
+$\mathbf{u}(k)\leftarrow M_{u}^{-1}\mathbf{u}(k),\quad\mathbf{y}(k)\leftarrow M_{y}^{-1}\mathbf{y}(k),$ (44)
+where $M_{u}=\text{diag}(m_{u_{1}},\dots,m_{u_{n_{u}}})$ and $M_{y}=\text{diag}(m_{y_{1}},\dots,m_{y_{n_{y}}})$.
+In this study, the scaling factors are chosen as the mean absolute values of the signals: $m_{u_{i}}=\text{mean}(|\mathbf{u}_{i}|)$ and $m_{y_{i}}=\text{mean}(|\mathbf{y}_{i}|)$.
+The computed optimal control input $\mathbf{u}_{N}^{*}(k)$ must be denormalized before being applied to the plant.
 Algorithm 1 Implementation of NTDPC
@@
 7: Solve the QP problem (Problem 4) and obtain $\mathbf{u}_{N}^{*}(k)$
-8:  Denormalize the signal $\mathbf{u}_{N}^{*}(k)$
-9:  Apply $u(k)=\mathbf{u}_{N_{[0:n_{u}-1]}}^{*}(k)$ to the system
+8:Extract optimal first input $u^{*}(k)$ from $u_{N}^{*}(k)$
+9:Denormalize $u^{*}(k)\leftarrow M_{u}u^{*}(k)$ and apply to system
 4 Stability Analysis
 This section establishes a rigorous stability analysis of the developed NTDPC approach.
-For both noise-free and noisy cases, we establish stability to the desired equilibrium.
-In the noise-free case, asymptotic stability ensures tracking of the reference.
-In the noisy case, input-to-state stability (ISS) guarantees robustness to bounded measurement noise.
-To facilitate the stability analysis, we define key concepts and notations as done in [24] for DeePC.
+We examine the system’s behavior in both noise-free and noisy scenarios.
+In the noise-free case, we prove asymptotic stability, ensuring exact tracking of the reference.
+In the noisy case, we demonstrate Input-to-State Stability (ISS), guaranteeing robustness to bounded measurement noise.
+To facilitate this analysis, we define the following key concepts and notations, adapted from the dissipativity framework for data-driven control [22].
 Definition 4.1 (Stage Cost).
-Assume $Q,R$ are block-diagonal:
-$\|\mathbf{y}_{N}(k)\|_{Q}^{2}$ $=\sum_{i=0}^{N-1}\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2},$
-$\|\mathbf{u}_{N}(k)\|_{R}^{2}$ $=\sum_{i=0}^{N-1}\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2},$ (52)
-where $Q_{i},R_{i}\succ 0$, and $r_{y}\in\mathbb{R}^{n_{y}}$, $r_{u}\in\mathbb{R}^{n_{u}}$ are the steady-state output and input, while $y^{*}(i|k),u^{*}(i|k)$ are the predicted output and input at step $i$ from time $k$.
-The stage cost is:
-$l(y^{*}(i|k),u^{*}(i|k))$ $=\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2}.$
-Thus, $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}$.
+Assume the weighting matrices $Q$ and $R$ are block-diagonal with blocks $Q_{i}\succ 0$ and $R_{i}\succ 0$. The stage cost $l(\cdot,\cdot)$ is defined as:
+$l(\mathbf{y},\mathbf{u})=\|\mathbf{y}-r_{y}\|_{Q_{i}}^{2}+\|\mathbf{u}-r_{u}\|_{R_{i}}^{2},$ (45)
+where $r_{y}\in\mathbb{R}^{n_{y}}$ and $r_{u}\in\mathbb{R}^{n_{u}}$ denote the steady-state output and input references, respectively.
+The total cost function over the prediction horizon $N$ is given by:
+$J^{*}(k)=\sum_{i=0}^{N-1}l(\mathbf{y}^{*}(i|k),\mathbf{u}^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2},$ (46)
+where $\mathbf{y}^{*}(i|k)$ and $\mathbf{u}^{*}(i|k)$ are the predicted output and input at step $i$ given information at time $k$.
 Definition 4.2 (Non-Minimal State).
-The non-minimal state is:
-$X_{\text{ini}}(k)$ $=\mathbf{z}_{\text{ini}}(k)=\text{col}(\mathbf{y}_{\text{ini}}(k),\mathbf{u}_{\text{ini}}(k))=\text{col}(y(k-T_{\text{ini}}),\dots,y(k-1),u(k-T_{\text{ini}}),\dots,u(k-1)),$
-with $\mathbf{y}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}},\mathbf{u}_{\text{ini}}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}}$. The error state is:
-$\tilde{X}_{\text{ini}}(k)$ $=\text{col}(\mathbf{y}_{\text{ini}}(k)-\mathbf{r}_{y},\mathbf{u}_{\text{ini}}(k)-\mathbf{r}_{u}),$
-where $\mathbf{r}_{y}=\text{col}(r_{y},\dots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\dots,r_{u})$.
+The non-minimal state vector $\mathbf{z}_{\text{ini}}(k)$ is defined as:
+$\mathbf{z}_{\text{ini}}(k)=\operatorname{col}(\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini}}(k)).$
+The error state $\tilde{\mathbf{z}}_{\text{ini}}(k)$ represents the deviation from equilibrium:
+$\tilde{\mathbf{z}}_{\text{ini}}(k)=\operatorname{col}(\mathbf{u}_{\text{ini}}(k)-\mathbf{r}_{u},\mathbf{y}_{\text{ini}}(k)-\mathbf{r}_{y}),$
+where $\mathbf{r}_{u}\in\mathbb{R}^{n_{u}T_{\text{ini}}}$ and $\mathbf{r}_{y}\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ are stacked reference vectors.
 Definition 4.3 (Class $\mathcal{K}$ and $\mathcal{KL}$ Functions).
-A function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\varphi(0)=0$.
-It is $\mathcal{K}_{\infty}$ if $\lim_{s\to\infty}\varphi(s)=\infty$.
-A function $\beta:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is $\mathcal{KL}$ if $\beta(\cdot,t)\in\mathcal{K}$ for fixed $t\geq 0$, and $\beta(s,t)\to 0$ as $t\to\infty$ for fixed $s>0$.
-The identity is $\text{id}(s)=s$.
+A continuous function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to class $\mathcal{K}$ if it is strictly increasing and $\varphi(0)=0$.
+It belongs to class $\mathcal{K}_{\infty}$ if it is also unbounded, i.e., $\lim_{s\to\infty}\varphi(s)=\infty$.
+A function $\beta:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to class $\mathcal{KL}$ if $\beta(\cdot,t)\in\mathcal{K}$ for each fixed $t\geq 0$, and $\lim_{t\to\infty}\beta(s,t)=0$ for each fixed $s>0$.
+The identity function is denoted by $\operatorname{id}(s)=s$.
 Definition 4.4 (Filtration).
-The filtration $\mathcal{F}_{k}$ is the sigma-algebra generated by the measurements $\{\mathbf{z}_{\text{ini},m}(t)\}_{t=0}^{k}$, representing the information available up to time $k$.
+The filtration $\mathcal{F}_{k}$ is the $\sigma$-algebra generated by the history of measurements $\{\mathbf{z}_{\text{ini},m}(t)\}_{t=0}^{k}$, representing the information available to the controller up to time $k$.
 4.1 Noise-Free Case
-We analyze the stability of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$), tracking references $(r_{y},r_{u})$.
+We first analyze the stability of the NTDPC scheme in the ideal noise-free scenario ($\mathbf{\zeta}_{\text{ini}}(k)=0$) while tracking constant admissible references $(r_{y},r_{u})$.
 Assumption 4.1 (Positive Definiteness).
-There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that for all $(y,u)\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$, where $\mathcal{Y}_{i}\subseteq\mathbb{R}^{n_{y}}$ and $\mathcal{U}_{i}\subseteq\mathbb{R}^{n_{u}}$ are the projections of $\mathcal{Y}$ and $\mathcal{U}$ onto the $i$-th prediction step:
-$\alpha_{1}(\|\text{col}(y-r_{y},u-r_{u})\|)\leq l(y,u)\leq\alpha_{2}(\|\text{col}(y-r_{y},u-r_{u})\|),$
-where $l(y,u)=\|y-r_{y}\|_{Q_{i}}^{2}+\|u-r_{u}\|_{R_{i}}^{2}$, and $(r_{y},r_{u})\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$ is a constant admissible equilibrium.
+There exist functions $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that for all admissible pairs $(\mathbf{y},\mathbf{u})$:
+$\alpha_{1}(\|\operatorname{col}(\mathbf{y}-r_{y},\mathbf{u}-r_{u})\|)\leq l(\mathbf{y},\mathbf{u})\leq\alpha_{2}(\|\operatorname{col}(\mathbf{y}-r_{y},\mathbf{u}-r_{u})\|).$ (47)
 Assumption 4.2 (Terminal Stabilizing Condition).
-For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<\text{id}$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
-$l(\hat{y}(N|k),\hat{u}(N|k))\leq(\text{id}-\rho)\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$
+For any admissible initial state $\mathbf{z}_{\text{ini}}(k)$, there exists a prediction horizon $N\geq T_{\text{ini}}$ and a terminal control law satisfying a contraction property.
+Specifically, there exists a function $\rho\in\mathcal{K}_{\infty}$ with $\rho<\operatorname{id}$ such that the terminal cost satisfies:
+$l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))\leq(\operatorname{id}-\rho)\circ l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (48)
+This condition ensures that the cost of the appended terminal state is strictly less than the cost of the oldest data point being discarded from the history window.
 Assumption 4.3 (Feasibility).
-Problem 4 is feasible for all $k\geq T_{\text{ini}}$ with $\mathbf{\zeta}_{\text{ini}}(k)=0$, and the optimal slack satisfies $\sigma_{y}^{*}(k)=0$.
-The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
+The NTDPC optimization (Problem 4) is recursively feasible for all $k\geq T_{\text{ini}}$.
+In the noise-free case, the optimal slack variable satisfies $\sigma_{y}^{*}(k)=0$.
+Furthermore, the equilibrium pair $(r_{y},r_{u})$ lies within the interior of the constraint sets $\mathcal{Y}\times\mathcal{U}$.
 Theorem 4.4.
-Under Assumptions 4.1, 4.2, and 4.3, with $\mathbf{\zeta}_{\text{ini}}(k)=0$ and a constant admissible equilibrium $(r_{y},r_{u})$, the closed-loop system under NTDPC is asymptotically stable, i.e., $y(k)\to r_{y}$, $u(k)\to r_{u}$ as $k\to\infty$.
+Under Assumptions 4.1, 4.2, and 4.3, consider the closed-loop system controlled by NTDPC in the absence of noise ($\mathbf{\zeta}_{\text{ini}}(k)=0$).
+The equilibrium $(r_{y},r_{u})$ is asymptotically stable, implying $\lim_{k\to\infty}\mathbf{y}(k)=r_{y}$ and $\lim_{k\to\infty}\mathbf{u}(k)=r_{u}$.
 Proof.
-Define the non-minimal state as in Definition 4.2:
-$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)).$
-The error state is $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$.
-Define the auxiliary function as $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ where the stage cost is $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ with $Q_{i},R_{i}\succ 0$.
-The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k)),$ since $\sigma_{y}^{*}(k)=0$ in the noise-free case (Assumption 4.3).
-Consider the storage function [24]:
-$V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k)).$
-To show positive definiteness of $V(\mathbf{z}_{\text{ini}}(k))$, using Assumption 4.1, we have:
-$l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
-so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|).$ Since $\tilde{X}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}})-r_{y},\ldots,u(k-1)-r_{u})$, there exists $\alpha_{1,W}\in\mathcal{K}_{\infty}$ such that $W(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|).$ Since $J^{*}(k)\geq 0$, there exists $\alpha_{1,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ For the upper bound:
-$l(y(k-i),u(k-i))\leq\alpha_{2}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
-so $W(\mathbf{z}_{\text{ini}}(k))\leq T_{\text{ini}}\alpha_{2}\left(\max_{i=1,\dots,T_{\text{ini}}}\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|\right).$ For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)$ (Lemma 3.5) depends on $\mathbf{z}_{\text{ini}}(k)=\tilde{X}_{\text{ini}}(k)+\text{col}(\mathbf{r}_{y},\mathbf{r}_{u})$.
-Bounded sets $\mathcal{Y},~{}\mathcal{U}$ (Assumption 4.3) and the predictor’s linearity ensure that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$.
-Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that $J^{*}(k)\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|).$ Combining, there exists $\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Hence, $\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Next, by defining the supply function as:
-$s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))=l(\hat{y}(N|k),\hat{u}(N|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})),$
-where $\hat{y}(N|k)$, $\hat{u}(N|k)$ satisfy Assumption 4.2. Construct a suboptimal solution at $k+1$:
-$\mathbf{u}_{s}(k+1)$ $=\text{col}(u^{*}(1|k),\ldots,u^{*}(N-1|k),\hat{u}(N|k))$
-$\mathbf{y}_{s}(k+1)$ $=\text{col}(y^{*}(1|k),\ldots,y^{*}(N-1|k),\hat{y}(N|k)),$ (53)
-with $\sigma_{y,s}(k+1)=0$.
-The suboptimal cost is as $J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k)).$ Since $J^{*}(k)=\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))$ and $J^{*}(k+1)\leq J_{s}(k+1)$:
-$J^{*}(k+1)\leq J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))$ (54)
-For the auxiliary function $W$:
-$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (55)
-Using (61) and (55), thus:
-$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))=[J^{*}(k+1)-J^{*}(k)]+[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))]$ (56)
-Then, we can rewrite the preceding equality in (56) as:
-$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))$ $\leq[J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))]$
-$\quad-J^{*}(k)+[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))]$
-$=s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (57)
-By Assumption 4.2, $s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\leq-\rho\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$ Since $l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\geq\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)$, we can rewrite the inequality (57) as:
-$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))\leq-\rho\circ\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)\leq-\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$
-Since $V$ is positive definite and radially unbounded with respect to $\tilde{X}_{\text{ini}}(k)$, and satisfies a strict dissipation inequality, standard Lyapunov results [25] imply $\tilde{X}_{\text{ini}}(k)\to 0$ as $k\to\infty$.
-Thus, $y(k)\to r_{y}$, $u(k)\to r_{u}$, proving asymptotic stability.
-Since $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ and $W(\mathbf{z}_{\text{ini}}(k))\geq 0$:
-$J^{*}(k)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|),$
-ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state. This completes the proof. ∎
+We construct a Lyapunov candidate using the optimal cost function augmented with a history storage function.
+Let the optimal cost at time $k$ be $J^{*}(k)$.
+Since $\sigma_{y}^{*}(k)=0$ (Assumption 4.3), $J^{*}(k)=\sum_{i=0}^{N-1}l(\mathbf{y}^{*}(i|k),\mathbf{u}^{*}(i|k))$.
+Define the auxiliary history function $W(\mathbf{z}_{\text{ini}}(k))$ as the accumulated cost of the past $T_{\text{ini}}$ steps:
+$W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(\mathbf{y}(k-i),\mathbf{u}(k-i)).$ (49)
+Consider the Lyapunov function candidate:
+$V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k)).$ (50)
+Using Assumption 4.1, the history function is bounded below by:
+$W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\operatorname{col}(\mathbf{y}(k-i)-r_{y},\mathbf{u}(k-i)-r_{u})\|).$ (51)
+Since the error state $\tilde{\mathbf{z}}_{\text{ini}}(k)$ is composed of these deviations, there exists $\alpha_{1,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\geq W(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)$.
+Similarly, using the upper bound $\alpha_{2}$ and the linearity of the predictor $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)$, the optimal cost $J^{*}(k)$ is bounded by the state magnitude.
+Thus, there exists $\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)$.
+At time $k+1$, we construct a feasible suboptimal input sequence $\mathbf{u}_{s}(k+1)$ by shifting the optimal sequence from time $k$ and appending the terminal law $\hat{\mathbf{u}}(N|k)$:
+$\mathbf{u}_{s}(k+1)=\operatorname{col}(\mathbf{u}^{*}(1|k),\ldots,\mathbf{u}^{*}(N-1|k),\hat{\mathbf{u}}(N|k)).$ (52)
+The corresponding cost is $J_{s}(k+1)=J^{*}(k)-l(\mathbf{y}^{*}(0|k),\mathbf{u}^{*}(0|k))+l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))$. By optimality, $J^{*}(k+1)\leq J_{s}(k+1)$.
+Now consider the evolution of the history function $W$:
+$W(\mathbf{z}_{\text{ini}}(k+1))$ $=\sum_{i=1}^{T_{\text{ini}}}l(\mathbf{y}(k+1-i),\mathbf{u}(k+1-i))$
+$=W(\mathbf{z}_{\text{ini}}(k))+l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (53)
+Note that in the noise-free prediction, $\mathbf{y}^{*}(0|k)=\mathbf{y}(k)$ and $\mathbf{u}^{*}(0|k)=\mathbf{u}(k)$.
+Computing the difference $\Delta V(k)=V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))$:
+$\Delta V(k)$ $\leq[J^{*}(k)-l(\mathbf{y}(k),\mathbf{u}(k))+l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))]$
+$\quad+[W(\mathbf{z}_{\text{ini}}(k))+l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}}))]$
+$\quad-[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))]$
+$=l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (54)
+Applying Assumption 4.2 (Terminal Stabilizing Condition):
+$\Delta V(k)\leq-\rho\circ l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}})).$ (55)
+Since $l(\cdot)$ is positive definite with respect to the deviation from equilibrium, this implies $\Delta V(k)$ is negative definite in terms of the error state.
+Standard Lyapunov arguments [23] then guarantee that $\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|\to 0$ as $k\to\infty$, confirming asymptotic stability.
+We assume that the system is stabilizable and that an admissible input sequence exists that maintains feasibility beyond the prediction horizon.
+4.2 Noisy Case: Robustness to Measurement Noise
+We now extend the analysis to the case where the measurements are corrupted by noise ($\mathbf{\zeta}_{\text{ini}}(k)\neq 0$).
+We prove that the NTDPC closed-loop system is Input-to-State Stable (ISS) with respect to the measurement noise $\mathbf{\zeta}_{\text{ini}}(k)$, utilizing the stochastic stability framework for MPC [23, 2].
+Assumption 4.5 (Bounded Noise Variance).
+The measurement noise is a martingale difference sequence with bounded variance. Specifically, there exists a constant $\bar{c}>0$ such that for all $k\geq 0$:
+$\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}\mid\mathcal{F}_{k}]\leq\bar{c},$ (56)
+where $\mathcal{F}_{k}$ denotes the filtration generated by measurements up to time $k$.
+Assumption 4.6.
+[Lipschitz Continuity of the Predictor] The predictor mapping is Lipschitz continuous with respect to the initial condition.
+There exists a constant $L_{P}>0$ such that for any two initial conditions $\mathbf{z}_{1},\mathbf{z}_{2}$:
+$\|\mathbf{P}_{1}(\mathbf{z}_{1}-\mathbf{z}_{2})\|\leq L_{P}\|\mathbf{z}_{1}-\mathbf{z}_{2}\|.$ (57)
+Consequently, the deviation in the slack variable due to noise is linearly bounded:
+$\|\sigma_{y,s}(k)\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k)\|,$ (58)
+where $\mathbf{z}_{\text{ini},m}(k)=\mathbf{z}_{\text{ini}}(k)+\operatorname{col}(0,\mathbf{\zeta}_{\text{ini}}(k))$.
+Assumption 4.7 (Stochastic Terminal Stabilizing Condition).
+For any admissible state, there exists a terminal control law $\hat{\mathbf{u}}(N|k)$ and a class $\mathcal{K}_{\infty}$ function $\rho<\operatorname{id}$ such that the expected terminal cost satisfies a contraction:
+$\mathbb{E}[l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))\mid\mathcal{F}_{k}]\leq(\operatorname{id}-\rho)\circ\mathbb{E}[l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}}))\mid\mathcal{F}_{k}].$ (59)
+Assumption 4.8.
+[Recursive Feasibility and Slack Bound] Problem 4 is feasible with probability at least $1-\epsilon$.
+Furthermore, provided the slack penalty weighting matrix $\Lambda_{y}$ is chosen sufficiently large relative to the noise covariance, the optimal slack variable $\sigma_{y}^{*}(k)$ satisfies a quadratic bound with respect to the noise realization:
+[…]
+for some constant $c_{\sigma}>0$.
 Remark 4.1.
-The bound $J^{*}(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ satisfies the requirement for a $\mathcal{K}_{\infty}$ cost bound, derived within the storage function analysis.
-4.2 Noisy Case: Robustness to Measurement Noise
-We analyze the stability of NTDPC under measurement noise ($\mathbf{\zeta}_{\text{ini}}(k)\neq 0$), tracking constant admissible references $(r_{y},r_{u})$.
-The system is shown to be input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$, following the framework in [2, 25].
-Assumption 4.5 (Bounded Noise Variance).
-There exists $\bar{c}>0$ such that $\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq\bar{c}$ for all $k$, where $\mathcal{F}_{k}$ is the filtration up to time $k$.
-Assumption 4.6 (Lipschitz Predictor).
-There exists $L_{P}>0$ such that:
-$\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k)-\mathbf{z}_{\text{ini}}(k))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k)\|,$
-where $L_{P}=\|\mathbf{P}_{1}\|_{2}$ is the spectral norm of $\mathbf{P}_{1}$, and $\mathbf{z}_{\text{ini},m}(k)=\mathbf{z}_{\text{ini}}(k)+\text{col}(0,\mathbf{\zeta}_{\text{ini}}(k))$.
-Assumption 4.7 (Terminal Stabilizing Condition - Noisy).
-For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<id$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
-$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(\text{id}-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$
-Assumption 4.8 (Feasibility - Noisy).
-Problem 4 is feasible with probability at least $1-\epsilon$, for small $\epsilon>0$. There exists $c_{\sigma}>0$ such that:
-$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}].$
-The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
-Remark 4.2.
-Assumption 4.7 is satisfied by selecting $\hat{u}(N|k)=r_{u}$ and leveraging the predictor’s robustness (Lemma 3.5).
-Assumption 4.8 is supported by the sensitivity index $I_{s}\leq 0.7$ (Fig.
-5), ensuring noise separability and feasibility.
+Assumption 4.7 is typically satisfied by selecting $\hat{\mathbf{u}}(N|k)$ as the steady-state reference $r_{u}$, utilizing the inherent robustness of the predictor established in Lemma 3.4.
+Assumption 4.8 is empirically supported by the sensitivity index analysis ($I_{s}\leq 0.7$), which ensures sufficient separability between signal and noise subspaces to maintain feasibility.
 Theorem 4.9.
-Under Assumptions 4.1, 4.5, 4.6, 4.7, and 4.8, with a constant admissible equilibrium $(r_{y},r_{u})$, NTDPC is input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$.
-There exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
-$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
+Under Assumptions 4.1, 4.5, 4.6, 4.7, and 4.8, the closed-loop system generated by NTDPC is Input-to-State Stable (ISS) with respect to the noise sequence $\mathbf{\zeta}_{\text{ini}}(k)$.
+Specifically, there exist functions $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ such that:
+$\mathbb{E}[\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{\mathbf{z}}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$ (61)
 Proof.
-Define the non-minimal state as in Definition 4.2:
-$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)),$
-and the error state $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$.
-The auxiliary function is $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ with stage cost $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ and $Q_{i},R_{i}\succ 0$.
-The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}.$ Define the stochastic Lyapunov function:
-$V(k)=\mathbb{E}[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$
-In order to study positive definiteness of $V(k)$, under Assumption 4.1, we have $l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$ so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|)\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|),$ for some $\alpha_{1,W}\in\mathcal{K}_{\infty}$.
-For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)+\sigma_{y}^{*}(k)$ and bounded sets $\mathcal{Y},\mathcal{U}$ (Assumption 4.8) imply that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$ and noise effects.
-By Assumption 4.8:
-$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}.$
-Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that:
-$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
-Combining, there exist $\alpha_{1,V},\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that:
-$\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
-As dissipation inequality, we construct a suboptimal solution at $k+1$ like (53) with $\sigma_{y,s}(k+1)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k+1)+\mathbf{P}_{2}\mathbf{u}_{s}(k+1)-\mathbf{y}_{s}(k+1)$, where $\hat{u}(N|k),\hat{y}(N|k)$ satisfy Assumption 4.7.
-The suboptimal cost is:
-$J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}.$
-By computing $\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]$ as follows:
-$\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (58)
-Since:
-$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (59)
-Then, by subtracting (58) and (59), we have:
-$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]$ $=-\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]-\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
-$\quad+\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]+\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
-By Assumption 4.7:
-$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(id-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (60)
-For the slack terms, by Assumption 4.6:
-$\|\sigma_{y,s}(k+1)\|\leq\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k+1)-\mathbf{z}_{\text{ini}}(k+1))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k+1)\|$ (61)
-so, by taking the expectation from both side of (61):
-$\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k+1)\|^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\bar{c}.$
-Similarly, $\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}$. Let $c_{\sigma,\text{max}}=\max(c_{\sigma},\lambda_{\max}(\Lambda_{y})L_{P}^{2})$. Then:
-$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$ (62)
-And for $W$:
-$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (63)
-By taking the expectation from both side of (63), thus:
-$\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]=\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (64)
-By combining (62) and (64):
-$V(k+1)-V(k)\leq\mathbb{E}[J_{s}(k+1)-J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$ (65)
-By substituting (60) in (65), yields:
-$V(k+1)-V(k)\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
-By Assumption 4.1, there exists $\alpha_{4}\in\mathcal{K}_{\infty}$ such that:
-$\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]\geq\alpha_{4}(\mathbb{E}[\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|])$ (66)
-Since $\tilde{X}_{\text{ini}}(k)$ includes these terms and using (66), there exists $\alpha_{4,V}\in\mathcal{K}_{\infty}$ such that:
-$V(k+1)-V(k)\leq-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
-The term $\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]$ is bounded by the predictor’s output and noise effects, contributing to the ISS gain.
-To show ISS property, By Assumption 4.5, $\bar{c}\leq c_{\zeta}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2}$, the dissipation inequality:
-$V(k+1)\leq V(k)-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\alpha_{5}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+c_{\zeta,\text{max}}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2},$
-where $\alpha_{5}\in\mathcal{K}_{\infty}$ bounds the control cost, and $c_{\zeta,\text{max}}=2c_{\sigma,\text{max}}c_{\zeta}$.
-Applying stochastic ISS results [25], there exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
-$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
-Since $V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}$ and $\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]\geq 0$:
-$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c},$
-ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state plus a noise-dependent term. Hence, the proof is completed. ∎
-Remark 4.3.
-The ISS property ensures bounded tracking errors under noise, as validated in Fig.
-2, where NTDPC maintains stable tracking of $(r_{y},r_{u})$ with $\sigma_{\zeta}^{2}=0.25I$.
-The sensitivity index $I_{s}$ (Fig.
-5) supports Assumption 4.8, aiding horizon selection.
+Consider the stochastic Lyapunov function candidate:
+$V(k)=\mathbb{E}[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))\mid\mathcal{F}_{k}].$ (62)
+From the noise-free analysis, $W(\mathbf{z}_{\text{ini}}(k))$ is lower-bounded by $\alpha_{1,W}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)$.
+The optimal cost $J^{*}(k)$ differs from the noise-free case by the slack term.
+Using Assumption 4.8 and the boundedness of the noise:
+$\mathbb{E}[J^{*}(k)\mid\mathcal{F}_{k}]\leq\alpha_{2,J}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$ (63)
+Thus, $V(k)$ satisfies the sandwich bounds:
+$\alpha_{1,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)\leq V(k)\leq\alpha_{2,V}(\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$ (64)
+We construct the standard suboptimal sequence $\mathbf{u}_{s}(k+1)$ by shifting the optimal input and appending the terminal law.
+The slack variable for this suboptimal solution, $\sigma_{y,s}(k+1)$, absorbs the prediction error due to noise.
+From Assumption 4.6, its expected cost is bounded:
+$\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}\mid\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\bar{c}.$ (65)
+Define $\bar{C}_{\text{noise}}=(c_{\sigma}+\lambda_{\max}(\Lambda_{y})L_{P}^{2})\bar{c}$. The expected difference in the Lyapunov function is:
+$\mathbb{E}[V(k+1)-V(k)\mid\mathcal{F}_{k}]$ $\leq\mathbb{E}\left[l(\hat{\mathbf{y}}(N|k),\hat{\mathbf{u}}(N|k))-l(\mathbf{y}(k-T_{\text{ini}}),\mathbf{u}(k-T_{\text{ini}}))\right$.
+$\quad\left.+l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}^{*}(0|k),\mathbf{u}^{*}(0|k))\mid\mathcal{F}_{k}\right]+\bar{C}_{\text{noise}}.$ (66)
+The first line is bounded by $\rho(\cdot)$ via Assumption 4.7.
+The crucial term is the difference in the second line: $\Delta_{\text{stage}}=l(\mathbf{y}(k),\mathbf{u}(k))-l(\mathbf{y}^{*}(0|k),\mathbf{u}^{*}(0|k))$.
+In the noise-free case, $\mathbf{y}(k)=\mathbf{y}^{*}(0|k)$, so this term vanishes.
+In the noisy case, the measured output $\mathbf{y}(k)$ differs from the predicted output $\mathbf{y}^{*}(0|k)$ due to the noise realization at time $k$.
+Since $l(\cdot)$ is locally Lipschitz with constant $L_{l}$ on the compact constraint set:
+$\mathbb{E}[\Delta_{\text{stage}}\mid\mathcal{F}_{k}]\leq L_{l}\mathbb{E}[\|\mathbf{y}(k)-\mathbf{y}^{*}(0|k)\|\mid\mathcal{F}_{k}]\leq L_{l}\mathbb{E}[\|\mathbf{\zeta}(k)\|\mid\mathcal{F}_{k}]\leq L_{l}\sqrt{\bar{c}}.$ (67)
+Combining these results:
+$\mathbb{E}[V(k+1)-V(k)\mid\mathcal{F}_{k}]\leq-\rho(\alpha_{4,V}(\mathbb{E}[\|\tilde{\mathbf{z}}_{\text{ini}}(k)\|]))+\bar{C}_{\text{total}},$ (68)
+where $\bar{C}_{\text{total}}$ aggregates all noise-dependent constants.
+This inequality confirms that the Lyapunov function decreases outside a bounded region determined by the noise level, which is the definition of ISS [23].
 5 Numerical Simulations
@@
 The inputs and outputs are constrained as follows:
-$\begin{split}&U:=\left\{u\in\mathbb{R}^{2}:\begin{bmatrix}-20\\ -20\end{bmatrix}\leq u\leq\begin{bmatrix}20\\ 20\end{bmatrix}\right\},\quad\\ &Y:=\left\{y\in\mathbb{R}^{2}:\begin{bmatrix}-25\\ -15\end{bmatrix}\leq y\leq\begin{bmatrix}25\\ 15\end{bmatrix}\right\}.\end{split}$ (68)
+[…]
 In this simulation, we set $T_{\text{ini}}=N=20$, data length for Hankel matrices $M=2500$.
@@
 The results are consistent with the asymptotic stability proven in Theorem 4.4, where $y(k)\to r_{y}$ and $u(k)\to r_{u}$ as $k\to\infty$ under Assumptions 4.1, 4.2, and 4.3.
-The top plot displays the primary output $y_{1}(k)$ for NTDPC (blue), SPC (red), and SMMPC (green) over time, with the dashed line representing the reference $r_{y}$; all methods converge to $r_{y}$, but NTDPC exhibits the fastest transient response and minimal overshoot, reflecting the effectiveness of the data-driven predictor and terminal condition.
-The second plot shows a secondary output $y_{2}(k)$, with similar convergence behavior, where NTDPC (blue) stabilizes more quickly than SPC (red) and SMMPC (green), supporting the robustness of the NTDPC framework to initial conditions encoded in $\mathbf{z}_{\text{ini}}(k)$.
-The third plot presents the primary input $u_{1}(k)$, where NTDPC (blue) adjusts rapidly from an initial value, stabilizing near $r_{u}$ (dashed line) with less oscillation compared to SPC (red) and SMMPC (green), consistent with the bounded input constraints in $\mathcal{U}$.
-Finally, the bottom plot depicts a secondary input $u_{2}(k)$, showing NTDPC (blue) achieving a steady state closer to $r_{u}$ with reduced variability, further validating the stability properties derived from the Lyapunov function $V(\mathbf{z}_{\text{ini}}(k))$.
+The top plot displays the primary output $y_{1}(k)$ for NTDPC (black), SPC (red), and SMMPC (green) over time, with the dashed line representing the reference $r_{y}$; all methods converge to $r_{y}$, but NTDPC exhibits the fastest transient response and minimal overshoot, reflecting the effectiveness of the data-driven predictor and terminal condition.
+The second plot shows a secondary output $y_{2}(k)$, with similar convergence behavior, where NTDPC (black) stabilizes more quickly than SPC (red) and SMMPC (green), supporting the robustness of the NTDPC framework to initial conditions encoded in $\mathbf{z}_{\text{ini}}(k)$.
+The third plot presents the primary input $u_{1}(k)$, where NTDPC (black) adjusts rapidly from an initial value, stabilizing near $r_{u}$ (dashed line) with less oscillation compared to SPC (red) and SMMPC (green), consistent with the bounded input constraints in $\mathcal{U}$.
+Finally, the bottom plot depicts a secondary input $u_{2}(k)$, showing NTDPC (black) achieving a steady state closer to $r_{u}$ with reduced variability, further validating the stability properties derived from the Lyapunov function $V(\mathbf{z}_{\text{ini}}(k))$.
 Overall, the convergence of outputs and inputs to their respective references demonstrates the asymptotic stability of NTDPC, as predicted by the storage function analysis $V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))$, and the superior performance of NTDPC over SPC and SMMPC highlights the advantage of the noise-tolerant predictor and optimization framework, particularly in the absence of noise.
@@
 In Fig.
-2, the simulation results demonstrate the performance of NTDPC in comparison to SPC and SMMPC under the noise effect, $\sigma_{\zeta}^{2}=0.25I$.
+3, the simulation results demonstrate the performance of NTDPC in comparison to SPC and SMMPC under the noise effect, $\sigma_{\zeta}^{2}=0.25I$.
 While SMMPC exhibits aggressive and unstable control actions leading to poor regulation of outputs, NTDPC and SPC show reasonable results, despite an initial overshoot and undershoot, ultimately achieves stable tracking of the reference for both outputs with damped oscillations and less aggressive control inputs that settle over time.
@@
 4 which reveals the superiority of NTDPC and SPC compared to SMMPC.
-It is worthy to note that there is a slight difference between NTDPC and SPC throughout the entire simulation, highlighting the effectiveness of the designed NTDPC.
+It is worth noting that there is a slight difference between NTDPC and SPC throughout the entire simulation, highlighting the effectiveness of the designed NTDPC.
 In summary, the figures provide a comprehensive comparison of NTDPC, SPC and SMMPC control strategies, highlighting their performance in terms of output and input trajectories and their respective variability.